### v3

# Introspectively Envy-Free and Efficient Allocation of Indivisible Mixed Manna

Abstract
The existence of allocations that are fair and efficient, simultaneously, is a central inquiry in fair division literature. A prominent result in discrete fair division shows that the complementary desiderata of fairness and efficiency can be achieved together when allocating indivisible items with nonnegative values; specifically, for indivisible goods and among agents with additive valuations, there always exists an allocation that is both envy-free up to one item ($\mathrm{EF}1$) and Pareto efficient ($\mathrm{PO}$). While a recent breakthrough extends the $\mathrm{EF}1$ and $\mathrm{PO}$ guarantee to indivisible chores (items with negative values), the question remains open for indivisible mixed manna, i.e., for indivisible items whose values can be positive, negative, or zero. The current work makes notable progress in positively resolving this central question.
For indivisible mixed manna and additive valuations, we establish the existence of allocations that are $\mathrm{PO}$ and introspectively envy-free up to one item ($\mathrm{IEF}1$). In an $\mathrm{IEF}1$ allocation, each agent can eliminate its envy towards all the other agents by either adding an item or removing an item from its own bundle. The notion of $\mathrm{IEF}1$ coincides with $\mathrm{EF}1$ for indivisible chores, and hence, our result generalizes the aforementioned existence guarantee for chores. Our techniques can also be adopted to obtain an alternative proof for the existence of $\mathrm{EF}1$ and $\mathrm{PO}$ allocations of indivisible goods. Hence, along with the $\mathrm{IEF}1$ and $\mathrm{PO}$ result for mixed manna, we provide a unified approach for establishing the $\mathrm{EF}1$ and $\mathrm{PO}$ guarantee for indivisible goods and indivisible chores, respectively. In addition, we utilize our result for indivisible items, along with limit arguments, to develop a distinct proof of the noted $\mathrm{EF}$ and $\mathrm{PO}$ guarantee for divisible mixed manna. Our work highlights an interesting application of the Knaster-Kuratowski-Mazurkiewicz (KKM) Theorem in discrete fair division and develops multiple, novel structural insights and algorithmic ideas.
1 Introduction
The existence of allocations that simultaneously achieve fairness and efficiency lies at the core of the fair division literature [BT96fair; moulinHandbookComputational16; moulinFairDivision03]. This question has been extensively studied over the past several decades, particularly in the context of goods – items with nonnegative values. Indeed, the classic result of Varian [varianEquityenvy74] establishes that, for divisible goods, a fair and efficient allocation always exists. The result lays a normative foundation for competitive markets with equal incomes: for divisible goods, a competitive equilibrium allocation (under equal incomes) is both envy-free—i.e., no agent values another’s bundle over their own [foleyResourceallocation67]—and Pareto efficient. In fact, such an allocation can be found by maximizing the Nash social welfare (the geometric mean of agents’ valuations); equivalently, by solving the Eisenberg-Gale convex program [eisenberg1959consensus].
While the compatibility of fairness and efficiency is well-studied for goods, it remains less resolved in the equally well-motivated settings of chores (negatively valued items) and, more generally, mixed manna (items whose valuations may vary in sign across agents). Notably, in contrast to the classic result for divisible goods [varianEquityenvy74], an analogous guarantee of fairness and efficiency for divisible mixed manna was established only relatively recently [Bogo2017]. The results in [Bogo2017] underscore that the mixed manna setting is significantly more complex than goods or even chores alone: For divisible mixed manna, unlike in the case of divisible goods, the set of competitive equilibria may be non-convex and even disconnected. Furthermore, although a competitive equilibrium allocation for divisible goods can be efficiently computed via a convex program, finding such an allocation in the divisible mixed manna setting is PPAD-hard [chaudhury2021competitive].
The current work addresses the compatibility between fairness and efficiency in the context of indivisible items, i.e., for resources and chores that cannot be fractionally assigned among the agents. For indivisible items as well, the separation between goods, chores, and mixed manna stands. For indivisible goods, a prominent result of Caragiannis et al. [caragiannis2019unreasonable] establishes that, under additive valuations, there always exists an allocation that is both Pareto efficient ($\mathrm{PO}$) and fair. Since an envy-free allocation is not guaranteed to exist for indivisible items, fairness in discrete fair division is captured through meaningful relaxations, most notably, envy-freeness up to one item ($\mathrm{EF}1$) [budishCombinatorialAssignment11]. An allocation of indivisible mixed manna is said to be $\mathrm{EF}1$ if, existing envy that any agent $i$ might have towards another agent $j$, can be eliminated by either the (hypothetical) removal of a single good from $j$’s bundle or the (hypothetical) removal of a chore from $i$’s bundle. [^1] When all the items are goods, as in [caragiannis2019unreasonable], this definition reduces to the requirement that any existing envy can be eliminated by removing a good from the other agent’s bundle. In particular, Caragiannis et al. [caragiannis2019unreasonable] showed that, under additive valuations, an allocation of indivisible goods that maximizes the Nash social welfare is both $\mathrm{EF}1$ and $\mathrm{PO}$. To compute $\mathrm{EF}1$ and $\mathrm{PO}$ allocations of indivisible goods, a pseudo-polynomial time algorithm—based on a Fisher market framework—was developed in [BKV2018]. Furthermore, when the number of agents is fixed, such allocations of goods can be computed efficiently [mahara2024].
In contrast to the established results for goods, the existence of $\mathrm{EF}1$ and $\mathrm{PO}$ allocations for indivisible chores—and, more generally, for mixed manna—remained open for nearly a decade. In fact, until the recent resolution by Mahara [mahara2025existence] for indivisible chores, the existence question was open even among four agents with additive valuations. Prior to [mahara2025existence], results for indivisible chores addressed special cases. For instance, [gargetalNewAlgorithms23] showed that $\mathrm{EF}1$ and $\mathrm{PO}$ allocations of indivisible chores exist among three agents. In addition, for chore division among $n$ agents, an (n1)(n-1) multiplicative approximation of $\mathrm{EF}1$ was considered in [GargConstantFactorSTOC25].
As mentioned, the work of Mahara [mahara2025existence] proves that an $\mathrm{EF}1$ and $\mathrm{PO}$ allocation of indivisible chores always exists among agents with additive valuations. Notably, however, the existence remains unresolved for indivisible mixed manna [^2] This is, in fact, explicitly posed as an open problem in [mahara2025existence]. – indivisible items whose values can be positive, negative, or zero.
[…]
The notion of $\mathrm{IEF}1$ coincides with $\mathrm{EF}1$ when all the indivisible items are chores, and hence, our result generalizes the existence guarantee for chores [mahara2025existence]. Our techniques can also be adopted to obtain alternate proof for the existence of $\mathrm{EF}1$ and $\mathrm{PO}$ allocations of indivisible goods. Hence, along with the $\mathrm{IEF}1$ and $\mathrm{PO}$ result for mixed manna, we provide a unified approach for establishing the $\mathrm{EF}1$ and $\mathrm{PO}$ guarantee for indivisible goods and indivisible chores, respectively.
Fairness and efficiency for indivisible mixed manna among $n$ agents is also considered in [barman2025fair]. The envy-freeness guarantee in [barman2025fair], however, requires a reallocation of up to (n1)(n-1) items. Both [mahara2025existence] and [barman2025fair] introduce novel applications of the Knaster-Kuratowski-Mazurkiewicz (KKM) theorem for the indivisible-items setting. Generalizing both of these results, [^3] An $\mathrm{IEF}1$ and $\mathrm{PO}$ allocation is also $\mathrm{EFR}$-(n1)(n-1) and $\mathrm{PO}$, where $\mathrm{EFR}$-(n1)(n-1) is the notion studied in [barman2025fair]. the current work advances the KKM-based framework for discrete fair division and develops new combinatorial ideas; Section 1.1 highlights our distinct technical contributions.
1.1 Our Results and Techniques
Our results are summarized next.
  • The current work establishes the existence of $\mathrm{IEF}1$ and $\mathrm{PO}$ allocations for indivisible mixed manna with additive valuations (Theorem 1).
  • Furthermore, our framework can be adapted to establish a fairness guarantee in terms of a chiral variant of $\mathrm{IEF}1$. In particular, one can also consider allocations that are extrospectively envy free up to one item ($\mathrm{XEF}1$) – each agent $i$ can eliminate its envy towards any other agent $j$ by either removing a good from $j$’s bundle or adding a chore to $j$’s bundle. Our KKM-based framework can be modified to prove that $\mathrm{XEF}1$ and $\mathrm{PO}$ allocations always exist (Theorem 28 in Appendix A).
    When all the indivisible items are goods, $\mathrm{XEF}1$ reduces to $\mathrm{EF}1$. Hence, our proof technique also leads to the $\mathrm{EF}1$ and $\mathrm{PO}$ guarantee for indivisible goods.
  • In addition, we establish that $\mathrm{IEF}1$ and fractionally Pareto efficient ($\mathrm{fPO}$) allocations always exist for indivisible mixed manna under additive valuations (Theorem 18). Note that an (integral) allocation is said to be fractionally Pareto efficient if it is not Pareto dominated even by fractional allocations of the items.
  • Finally, we prove that our $\mathrm{IEF}1$ and $\mathrm{fPO}$ result (Theorem 18), along with limit arguments, imply the existence of envy-free ($\mathrm{EF}$) and $\mathrm{fPO}$ allocations of divisible mixed manna (Theorem 20). This result for the divisible case provides an alternative proof of the existential guarantee in [Bogo2017] for additive valuations.
The novel technical ideas and algorithmic techniques developed in this work are outlined next.
For indivisible goods, $\mathrm{EF}1$ plus $\mathrm{PO}$ guarantees were obtained by either maximizing Nash social welfare [caragiannis2019unreasonable] or via the Fisher market based approach [BKV2018; mahara2025existence; garg2023computing]. It was shown in [eckart2024fairness] that there does not exist a welfare function (analogous to Nash social welfare) that guarantees $\mathrm{EF}1$ for chores (and, hence, for mixed manna). In addition, the market based approach does not directly extend to the mixed manna setting.
A successful alternative approach, which extends beyond the goods setting, was identified recently in [barman2025fair] and [mahara2025existence]. These results are based on novel applications of the Knaster-Kuratowski-Mazurkiewicz (KKM) theorem for the indivisible-items setting along with combinatorial arguments. However, the $\mathrm{EF}1$ and $\mathrm{PO}$ guarantee in [mahara2025existence] applies specifically to indivisible chores, with the mixed manna case explicitly stated as an open problem. The envy-freeness guarantee in [barman2025fair] holds for mixed manna, however, it requires a reallocation of up to (n1)(n-1) items among the $n$ agents.
Establishing the $\mathrm{IEF}1$ and $\mathrm{PO}$ guarantee for mixed manna requires new ideas on both the KKM front and in the subsequent combinatorial arguments. We highlight these below.
Invoking the KKM Theorem. A key difficulty that arises in the mixed manna setting is the presence of items that are valued zero by some agents and positively or negatively by others. In the setting with only chores [mahara2025existence], such items can be preprocessed: one can a priori remove such an item from consideration, identify a fair and efficient allocation in the remaining instance, and then subsequently assign the item to any agent that values it at zero. This fix maintains the fairness and the efficiency guarantee in the all-chores case. The KKM theorem in [mahara2025existence] is essentially used to find a (component-wise) positive weight vector for the agents. The final, fair and efficient allocation, is a weighted welfare-maximizing allocation with respect to these weights. In the presence of such zero-valued items, the covering condition required to invoke the KKM theorem breaks down. Moreover, the following example shows that, in the absence of an auxiliary fix, no positive weights can result in an allocation that satisfies a desired fairness requirement: Consider an instance with two agents and $m$ items. Agent $1$ values each item at $1$ and agent $2$ values each item at 0. Under any positive weights, $w_{1},w_{2}>0$, any weighted welfare-maximizing allocation assigns all the items to agent $1$. That is, under any such efficient allocation, agent $2$ receives an empty bundle. Hence, we get that agent $2$ can never be envy-free when considering prices (dual variables), i.e., agent $2$ can never be price envy-free (contrary to the desired property in Theorem 10).
To handle the issue with the zero-valued items, we construct a non-degenerate instance and introduce an auxiliary item with an identical, positive value for all the agents. The presence of this auxiliary good ensures that each agent, under suitable weights, can always receive at least one item, particularly the auxiliary good. This is a subtle, but key, step towards ensuring that, the covering condition for the KKM theorem (Lemma 9) holds, even in the mixed manna setting. That is, even for mixed manna, we are able to invoke the KKM Theorem (Theorem 7) for establishing the existence of a welfare-maximizing (efficient) and fair allocation (Theorem 10).
In our $\mathrm{XEF}1$ and $\mathrm{PO}$ result (Theorem 28), the definition of the sets ($\overline{C}_{i}$s) involved in the KKM application is also novel.
Furthermore, in contrast to the current work and [mahara2025existence], the definition of the KKM sets in [barman2025fair] does not involve prices (dual variables) and price envy-freeness. Hence, our arguments to invoke KKM differ from the ones in [barman2025fair].
We also additively perturb the non-zero values in the given fair division instance to construct a non-degenerate one. These perturbations provide us with crucial non-degeneracy properties (Lemmas 2 and 3). The construction of the non-degenerate instance relies on a collection of interdependent parameters that control the magnitude of the perturbations ($\varepsilon$), the value of the auxiliary item ($\lambda$), weight shifts for the agents ($\eta$), and other quantities (e.g., $\omega$). These parameters must be carefully defined, since the proofs delicately depend on their values and interrelationships. Hence, one of the technical contributions of this work is the construction of the non-degenerate instance with an auxiliary item (and carefully selected parameters) that addresses the challenges posed by zero-valued items and enables the application of the KKM theorem.
Combinatorial Insights and the Augmenting Tree Algorithm. Having identified, via the KKM theorem, a weight vector ($w^{*}$) and an associated collection of weighted welfare-maximizing allocations that are price envy-free for the $n$ agents, we proceed to constructively prove the existence of the desired fair and efficient allocation. [mahara2025existence] follows a similar high-level approach, albeit without the above-mentioned complexities of mixed manna. However, the subsequent argument and the proofs in [mahara2025existence] require all items to be chores, [^4] In particular, the proofs in Section 3.1 [mahara2025existence] crucially require positive prices. and hence do not extend to mixed manna. To address this, we establish new structural insights and develop a novel method to consolidate the collection of $n$ respective price envy-free allocations into a single fair allocation.
The consolidation is achieved via our augmenting tree algorithm, which iteratively reassigns items through a traversal of a tree comprising agents and items (i.e., the price graph). The guarantee that these tree augmentations can be successfully performed is based on a key structural lemma (Lemma 14). Here, we quantify progress using a potential defined in terms of a min-max price threshold. This threshold plays a crucial role in both the design and analysis of the algorithm. We also note that the analysis of our augmenting tree algorithm—in particular, its termination—is comparatively simpler than that in [mahara2025existence]. Hence, when considered specifically for the all-chores case, it provides a more direct proof of the existence of $\mathrm{EF}1$ and $\mathrm{PO}$ allocations.
1.2 Additional Related Work
$\mathrm{EF}1$ for Indivisible Goods, Chores, and Mixed Manna. The following prior works address $\mathrm{EF}1$ division of indivisible items, without the $\mathrm{PO}$ desideratum. For goods and under monotone valuations, an $\mathrm{EF}1$ allocation is guaranteed to exist and can be computed efficiently via the envy-cycle elimination algorithm [LMM04]. Furthermore, under additive valuations, the round-robin algorithm finds $\mathrm{EF}1$ allocations for both goods-only and chores-only settings.
[azizetalFairallocation22] established the existence and efficient computation of $\mathrm{EF}1$ allocations of mixed manna with additive valuations. [BSV21] showed that $\mathrm{EF}1$ allocation for chores with monotone (decreasing) valuations can be efficiently computed via the elimination of top-trading envy cycles.
$\mathrm{EF}1$ and $\mathrm{PO}$ for Mixed Manna. For the specific case of two agents, the existence of $\mathrm{EF}1$ and $\mathrm{PO}$ allocations of mixed manna was established in [azizetalFairallocation22]. This guarantee is also known particularly for identical or ternary valuations [aleksandrovwalshTwoAlgorithms20]. Also, for two agents and under category constraints, [SHS23] considered the relaxation $\mathrm{EF}{[1,1]}$ and showed that mixed-manna allocations that are $\mathrm{EF}{[1,1]}$ and $\mathrm{PO}$ exist and can be computed efficiently. Category constraints were further studied by Igarashi and Meunier [IM25], who established envy-freeness under n(n1)n(n-1) reallocations. For restricted additive valuations, in particular, [livanos2022almost] obtained fairness and efficiency guarantees for mixed manna. Overall, prior to the current work, efficiency and up-to-one-item envy-freeness guarantees for mixed manna were not known in the general setting of arbitrarily many agents with additive valuations.
$\mathrm{EF}1$ and $\mathrm{PO}$ for Chores. Before [mahara2025existence], $\mathrm{EF}1$ and $\mathrm{PO}$ results for indivisible chores were known only in special cases. $\mathrm{EF}1$ and $\mathrm{PO}$ chore allocations, with additive valuations, were known to exist for two agents [azizetalFairallocation22], for three agents [gargetalNewAlgorithms23], and for three types of agents [gargetalWeightedEF124].
Previous works have also established fairness and efficiency guarantees for restricted classes of valuations [ebadianetalHowfairly22; gargetalFairEfficientAAAI22; wuetalWeightedEF123; GargConstantFactorSTOC25; azizetalFairallocation23].
$\mathrm{Prop}1$ and $\mathrm{PO}$ for Mixed Manna. Share-based fairness criteria have also been studied for indivisible items [amanatidis2023fair]. In particular, an allocation of mixed manna is said to be proportional up to one item ($\mathrm{Prop}1$) if each agent can, by adding a good or removing a chore from its bundle, ensure a value at least as much as its proportional share, i.e., at least $1/n$ times its value for the grand bundle [azizetalFairallocation22]. Note that, in $\mathrm{Prop}1$ (as in $\mathrm{IEF}1$), items are added or removed from an agent’s own bundle. Moreover, under additive values, an $\mathrm{IEF}1$ allocation of mixed manna is also $\mathrm{Prop}1$. Therefore, our $\mathrm{IEF}1$ and $\mathrm{PO}$ result recovers the existential guarantee for $\mathrm{Prop}1$ and $\mathrm{PO}$ mixed-manna allocations, established in [aziz2020polynomial]; see also [branzei2024algorithms; barman2019proximity].
2 Preliminaries
Fair Division Instance. A fair division instance $\mathcal{I}$ is denoted by a triple $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$; here, $m\in\mathbb{Z}_{+}$ indivisible items need to be partitioned among $n\in\mathbb{Z}_{+}$ agents with individual valuations $v_{i}$. The set of agents and items will be denoted by $n]\coloneqq\{1,2,\ldots,n\}$ and $m]\coloneqq\{1,2,\ldots,m\}$, respectively. We will, throughout, write $v_{i}(t)\in\mathbb{R}$ to denote the value of each item $t\in[m]$ for agents $i\in[n]$. Furthermore, this work addresses agents with additive valuations, i.e., for each agent $i\in[n]$ and for any set of items $S\subseteq[m]$, the value $v_{i}(S)\coloneqq\sum_{t\in S}v_{i}(t)$. By convention, we have $v_{i}(\emptyset)=0$ for each agent $i\in[n]$. We obtain results for fair division instances with mixed manna, i.e., the values, $v_{i}(t)$, can be positive, negative, or zero.
For any subset $X\subseteq[m]$ and item $t\in[m]$, we will write $X+t$, as a shorthand, to denote the set $X\cup\{t\}$. Also, write XtX-t to denote $X\setminus\{t\}$. For any two subsets $X,Y\subseteq[m]$, we will use $X\triangle Y$ to denote the symmetric set difference between them, $X\triangle Y\coloneqq(X\setminus Y)\cup(Y\setminus X)$. Note that, here, if $Y$ is a singleton, then $\triangle$ toggles items in and out of $X$. Specifically, if $t\in X$, then it holds that $X\triangle\{t\}=X-t$. Otherwise, for $t\notin X$, we have $X\triangle\{t\}=X+t$.
Allocation. An allocation $\mathcal{A}=(A_{1},A_{2},\ldots,A_{n})$ is a partition of the $m$ items among the $n$ agents; specifically, for the item subsets $A_{i}\subseteq[m]$ we have $A_{i}\cap A_{j}=\emptyset$, for all $i\neq j$, and $\cup_{i=1}^{n}A_{i}=[m]$. Here, the subset $A_{i}$ is assigned to agent $i$ and is referred to as agent $i$’s bundle. Furthermore, a partial allocation $(A^{\prime}_{1},A^{\prime}_{2},\ldots,A^{\prime}_{n})$ corresponds to a collection of pairwise-disjoint subsets ($A^{\prime}_{i}\cap A^{\prime}_{j}=\emptyset$ for all $i\neq j$) wherein some of the items may remain unassigned, $\cup_{i=1}^{n}A^{\prime}_{i}\subsetneq[m]$. Complementarily and to reinforce that all the items are assigned in an allocation, we will use the term complete allocation. The set of all complete allocations of the items, $[m]$, among the $n$ agents will be denoted by $\Pi_{n}([m])$.
A fractional allocation $\mathcal{X}=(X_{1},\ldots,X_{n})$ corresponds to a fractional assignment of the items; here, $X_{i}=(X_{i,1},X_{i,2},\ldots,X_{i,m})$ with each $X_{i,t}\in[0,1]$ denoting the fraction of item $t$ assigned to agent $i$. In a fractional allocation $\mathcal{X}=(X_{1},\ldots,X_{n})$ we have $\sum_{i=1}^{n}X_{i,t}=1$, for each item $t\in[m]$, and each agent $i$ receives a value $v_{i}(X_{i})\coloneqq\sum_{t=1}^{m}X_{i,t}\ v_{i}(t)$. Indeed, if $X_{i,t}\in\{0,1\}$ for each $i$ and $t$, then $\mathcal{X}$ is an (integral) allocation.
Fairness. Next, we define the fairness notions addressed in this work. A central fairness criterion in fair division is that of envy-freeness. Specifically, an allocation $\mathcal{S}=(S_{1},\ldots,S_{n})$ is said to be envy-free ($\mathrm{EF}$) if each agent $i\in[n]$ values its bundle at least as much as that of any other agent, $v_{i}(S_{i})\geq v_{i}(S_{j})$ for all $j\in[n]$. For indivisible items, an envy-free allocation is not guaranteed to exist, motivating the study of relaxations of envy-freeness in discrete fair division. A well-studied relaxation is envy-freeness up to one item ($\mathrm{EF}1$). Formally, an allocation $\mathcal{B}=(B_{1},\ldots,B_{n})$ of mixed manna is said to be $\mathrm{EF}1$ if, for every pair of agents $i,j\in[n]$, either $i$ does not envy $j$, or there exists an items $t\in B_{i}\cup B_{j}$ such that $v_{i}(B_{i}-t)\geq v_{i}(B_{j}-t)$. That is, existing envy from agent $i$ towards agent $j$ can be eliminated by either removing a chore from $i$’s bundle or removing a good from $j$’s bundle.
This work resolves the fairness and efficiency question for mixed manna, considering a close variant of the standard $\mathrm{EF}1$ criterion. In particular, we deem an allocation to be fair—introspectively envy-free up to one item ($\mathrm{IEF}1$)—if existing envy from agent $i$ towards agent $j$ can be resolved by removing a chore or adding a good to $i$’s own bundle. Hence, unlike $\mathrm{EF}1$, in $\mathrm{IEF}1$ the hypothetical inclusion or exclusion of an item $t$, aimed at eliminating $i$’s envy, only concerns $i$’s bundle, i.e., the notion is introspective. [^5] Recall that $X\triangle\{t\}=X-t$, if $t\in X$. Otherwise, for $t\notin X$, we have $X\triangle\{t\}=X+t$.
Definition 1 ($\mathrm{IEF}1$).
In an instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$, an allocation $\mathcal{A}=(A_{1},\ldots,A_{n})\in\Pi_{n}([m])$ is said to be introspectively envy-free up to one item ($\mathrm{IEF}1$) if for each agent $i\in[n]$ there exists a subset $S\subseteq[m]$ of size at most one ($|S|\leq 1$) such that $v_{i}(A_{i}\triangle S)\geq\max\limits_{1\leq j\leq n}\ v_{i}(A_{j})$.
Note that in fair division instances with only chores (i.e., with $v_{i}(t)<0$ for all agents $i$ and items $t$) an allocation is $\mathrm{IEF}1$ iff it is $\mathrm{EF}1$. In the case of indivisible goods (i.e., when $v_{i}(t)\geq 0$ for all agents $i$ and items $t$), $\mathrm{EF}1$ implies $\mathrm{IEF}1$: Starting with an $\mathrm{EF}1$ allocation $(A_{1},\ldots,A_{n})$ of goods, to satisfy Definition 1 for an envious agent $i$, we can set $S=\{t^{*}\}$, where good $t^{*}\in\operatorname*{arg\,max}_{t\in[m]\setminus A_{i}}\ v_{i}(t)$.
Efficiency and Social Welfare. In a fair division instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$, a (possibly fractional) allocation $\mathcal{X}=(X_{1},\ldots,X_{n})$ is said to Pareto dominate another allocation $\mathcal{Y}=(Y_{1},\ldots,Y_{n})$ if $v_{i}(X_{i})\geq v_{i}(Y_{i})$, for all $i\in[n]$, and this inequality is strict for at least one agent.
Definition 2 ($\mathrm{PO}$).
In an instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$, an allocation $\mathcal{A}\in\Pi_{n}([m])$ is said to be Pareto efficient ($\mathrm{PO}$) if no other allocation $\mathcal{B}\in\Pi_{n}([m])$ Pareto dominates it.
Definition 3 ($\mathrm{fPO}$).
In an instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$, an allocation $\mathcal{A}\in\Pi_{n}([m])$ is said to be fractionally Pareto efficient ($\mathrm{fPO}$) if no fractional allocation $\mathcal{X}\in[0,1]^{n\times m}$ Pareto dominates it.
The social welfare of any allocation $\mathcal{A}=(A_{1},\ldots,A_{n})$ is the sum of agents’ values under it, $\sum_{i=1}^{n}v_{i}(A_{i})$.
We will use $\Delta_{n-1}$ to denote the (n1)(n-1)-dimensional simplex, $\Delta_{n-1}\coloneqq\{w\in\mathbb{R}^{n}\mid\sum_{i=1}^{n}w_{i}=1\text{ and }w_{i}\geq 0\text{ for all }i\in[n]\}$. For any vector $w\in\mathbb{R}_{\geq 0}^{n}$, write $\mathrm{supp}(w)$ to denote the support (positive components) of vector $w$, i.e., $\mathrm{supp}(w)\coloneqq\{j\in[n]\mid w_{j}>0\}$.
3 Main Result: $\mathrm{IEF}1$ and $\mathrm{PO}$ allocations for Mixed Manna
The following theorem provides our main result, the guaranteed existence of a fair and efficient allocation of indivisible mixed manna.
Theorem 1.
Every fair division instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$, with indivisible mixed manna and additive valuations, admits an allocation that is both introspectively envy-free up to one item ($\mathrm{IEF}1$) and Pareto efficient ($\mathrm{PO}$).
Proof Structure.
The proof of Theorem 1 is developed in the following three sections.
First, Section 4 establishes that it suffices to focus on non-degenerate instances. In particular, starting from the given fair division instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$, we construct (in Section 4.1) a related instance $\widehat{\mathcal{I}}$. Lemmas 2 and 3 (in Section 4.1) provide the non-degeneracy properties of the constructed instance $\widehat{\mathcal{I}}$. Subsequently, Lemmas 4 and 5 (in Section 4.2) show that existence of a fair and efficient (in particular, a welfare-maximizing) allocation in $\widehat{\mathcal{I}}$ implies the desired existence of an $\mathrm{IEF}1$ and $\mathrm{PO}$ allocation in the given instance $\mathcal{I}$. Hence, Theorem 1 reduces to establishing fairness and efficiency in $\widehat{\mathcal{I}}$ (see Theorem 6).
Subsequently, Section 5 sets up the Knaster-Kuratowski-Mazurkiewicz (KKM) Theorem towards establishing the desired fairness and efficiency guarantee in $\widehat{\mathcal{I}}$. Here, we prove (via the KKM Theorem) the existence of a weight vector $w^{*}\in\Delta_{n-1}$ with the property that, for each agent $i\in[n]$, we have a weighted welfare-maximizing allocation $\mathcal{A}^{(i)}$ that is fair (in a price sense) for $i$; see Theorem 10.
Finally, in Section 6, we start with the weight vector $w^{*}$ and the corresponding welfare-maximizing allocations $\mathcal{A}^{(1)},\mathcal{A}^{(2)},\ldots,\mathcal{A}^{(n)}$ that are price envy-free for the $n$ agents, respectively. We establish that the existence of such a collection of allocations implies the existence of a single allocation $\mathcal{A}^{*}$ that is $\mathrm{IEF}1$ and weighted welfare-maximizing in $\widehat{\mathcal{I}}$; this implication is obtained via a novel Augmenting Tree Method (Theorem 13). Therefore, using the above-mentioned sufficiency of fairness and efficiency in $\widehat{\mathcal{I}}$, we overall obtain our main result, Theorem 1.
4 Sufficiency of Non-Degenerate Instances
Item Types.
Without loss of generality, we will solely focus on instances $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$ wherein all the items in $[m]$ are of three types: goods, chores, or non-negatively valued. An item $t\in[m]$ is said to be a good if $v_{i}(t)>0$ for all agents $i\in[n]$. An item $t\in[m]$ is a chore if $v_{i}(t)<0$ for all agents $i\in[n]$. Further, an item is non-negatively valued if $v_{i}(t)\geq 0$ for all agents $i\in[n]$ and we have $v_{a}(t)>0$ for some agent $a\in[n]$ along with $v_{b}(t)=0$ for some $b\in[n]$. This assumption that items are exclusively of these three types holds without loss of generality, given that we identify Pareto efficient (and fair) allocations. In particular, if in a given instance there exists an item $t$ that is positively valued by some agent and negatively valued by another agent, then we can set the value of $t$ to be zero for all agents that have a negative (or zero) value for it, and keep the values unchanged for agents that positively value $t$, i.e., transform $t$ into a non-negatively valued item. Note that, under the transformed valuation, in any Pareto efficient allocation, item $t$ must be assigned to an agent that positively valued it. This observation implies that any $\mathrm{PO}$ and $\mathrm{IEF}1$ allocation $\mathcal{A}$ under the transformed valuation is efficient and fair with respect to the given valuations as well. Furthermore, we can set aside any item $t^{\prime}$ with the property that $v_{i}(t^{\prime})\leq 0$, for all agents $i$, and $v_{a}(t^{\prime})=0$, for some agent $a$. Here, if $\mathcal{A}$ is a $\mathrm{PO}$ and $\mathrm{IEF}1$ allocation with $t^{\prime}$ removed from consideration, then subsequently assigning $t^{\prime}$ to any agent $a$ with $v_{a}(t^{\prime})=0$ maintains fairness and efficiency.
Starting from the given instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$, we will first construct a non-degenerate instance $\widehat{\mathcal{I}}=\langle[n],[\widehat{m}],\{\widehat{v}_{i}\}_{i=1}^{n}\rangle$, which will have certain desirable properties. We will show that the existence of a fair and efficient allocation in $\widehat{\mathcal{I}}$ implies the existence of an $\mathrm{IEF}1$ and $\mathrm{PO}$ allocation in $\mathcal{I}$. Hence, for the purpose of proving Theorem 1, we will address instance $\widehat{\mathcal{I}}$.
We begin by describing the construction of the instance $\widehat{\mathcal{I}}$ (in Section 4.1). We will then detail the (non-degeneracy) properties satisfied by $\widehat{\mathcal{I}}$ (in Lemmas 2 and 3), and show that for proving Theorem 1 it suffices to establish a fairness and efficiency guarantee in $\widehat{\mathcal{I}}$ (Lemmas 4 and 5).
4.1 Constructing Non-degenerate Instance $\widehat{\mathcal{I}}$
To construct $\widehat{\mathcal{I}}=\langle[n],[\widehat{m}],\{\widehat{v}_{i}\}_{i=1}^{m}\rangle$ from the given the instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$, we first define the following parameters
$\lambda\coloneqq\min_{i\in[n],S,T\subseteq[m]}\big\{v_{i}(S)-v_{i}(T)\mid v_{i}(S)-v_{i}(T)>0\big\},$
$\omega\coloneqq\min_{\mathcal{A},\mathcal{B}\in\Pi_{n}([m])}\left\{\sum_{i=1}^{n}v_{i}(A_{i})-\sum_{i=1}^{n}v_{i}(B_{i})\mid\sum_{i=1}^{n}v_{i}(A_{i})-\sum_{i=1}^{n}v_{i}(B_{i})>0\right\}.$
Note that $\lambda$ represents the minimum possible envy in the given instance $\mathcal{I}$ and $\omega$ is the minimum gap between social welfares. Both $\lambda$ and $\omega$ are positive. We further fix parameter $\varepsilon>0$ to satisfy
$\varepsilon<\frac{\lambda\omega}{16m^{2}n\ \max\limits_{i\in[n],t\in[m]}|v_{i}(t)|}$ (1)
All the terms in the right-hand-side of equation (1) are positive and, hence, we can fix an $\varepsilon>0$ that satisfies this strict inequality. In addition, note that $\omega\leq 2m\max\limits_{i\in[n],t\in[m]}|v_{i}(t)|$ – this bound gives us
$\varepsilon<\frac{\lambda}{8m}$ (2)
For the instance $\widehat{\mathcal{I}}=\langle[n],[\widehat{m}],\{\widehat{v}_{i}\}_{i=1}^{m}\rangle$, we set $\widehat{m}\coloneqq m+1$, i.e., we add a new item to the existing set $[m]$. For the new item, $(m+1)$, in the instance $\widehat{\mathcal{I}}$ we set the value $\widehat{v}_{i}(m+1)=\lambda/2$ for each agent $i\in[n]$. Furthermore, for agents $i\in[n]$ and items $t\in[m]$ with nonzero $v_{i}(t)$, we additively perturb this given value to obtain $\widehat{v}_{i}(t)$ in the instance $\widehat{\mathcal{I}}$. Specifically, for each $i$ and $t$, with $v_{i}(t)\neq 0$, independently and uniformly at random draw $\varepsilon_{i,t}\sim\mathrm{Uniform}([0,\varepsilon])$ and set $\widehat{v}_{i}(t)=v_{i}(t)-\varepsilon_{i,t}$. We preserve the zero values: if for any agent $a$ and item $s$, we have $v_{a}(s)=0$, then set $\widehat{v}_{a}(s)=0$. As in the given instance $\mathcal{I}$, the agents’ valuations, $\widehat{v}_{i}(\cdot)$, in $\widehat{\mathcal{I}}$ are additive. This completes the construction of $\widehat{\mathcal{I}}=\langle[n],[\widehat{m}],\{\widehat{v}_{i}\}_{i=1}^{n}\rangle$.
The proposition below notes that the sign of each value $v_{i}(t)$ is preserved by $\widehat{v}_{i}(t)$ and, in the two instances, the maximum absolute values across the items are comparable.
Proposition 1.
For any agent $i\in[n]$ and item $t$, if the given value $v_{i}(t)>0$, then $\widehat{v}_{i}(t)$ is positive as well. Similarly, if $v_{a}(s)=0$, for an agent $a$ and item $s$, then $\widehat{v}_{a}(s)=0$. In addition, for each agent $j$ and item $t^{\prime}$ with $v_{j}(t^{\prime})<0$, it holds that $\widehat{v}_{j}(t^{\prime})<0$. Furthermore,
$\max_{\begin{subarray}{c}i\in[n]\\ t\in[m]\end{subarray}}\ |v_{i}(t)|\ \geq\ \frac{1}{2}\max_{\begin{subarray}{c}i\in[n]\\ t\in[\widehat{m}]\end{subarray}}\ |\widehat{v}_{i}(t)|.$
Proof.
If $v_{a}(s)=0$, then, by construction, we have $\widehat{v}_{a}(s)=0$. Next, note that if $v_{j}(t^{\prime})<0$, then we set $\widehat{v}_{j}(t^{\prime})=v_{j}(t^{\prime})-\varepsilon_{j,t^{\prime}}$ with $\varepsilon_{j,t^{\prime}}\geq 0$. Hence, $v_{j}(t^{\prime})<0$ implies $\widehat{v}_{j}(t^{\prime})<0$.
In addition, for each value $v_{i}(t)>0$, it holds that $v_{i}(t)=v_{i}(t)-v_{i}(\emptyset)\geq\lambda$; the last inequality follows from the definition of $\lambda$. Now, using the fact $\varepsilon_{i,t}\leq\varepsilon<\frac{\lambda}{8m}$ (see equation (2)), we obtain $\widehat{v}_{i}(t)=v_{i}(t)-\varepsilon_{i,t}>\lambda-\frac{\lambda}{8m}>0$. Therefore, each positive value, $v_{i}(t)>0$, continues to be positive under the constructed valuation, $\widehat{v}_{i}(t)>0$.
We next relate the maximum absolute values across the items in the two instances $\mathcal{I}$ and $\widehat{\mathcal{I}}$. In the definition of $\lambda$, the minimization includes considering one of the sets to be a singleton and the other to be empty. Hence, for any nonzero value $v_{i}(t)$ it holds that $|v_{i}(t)|\geq\lambda$. Therefore, for any $i$ and $t$ with nonzero $v_{i}(t)$, we have
$|\widehat{v}_{i}(t)|=|v_{i}(t)-\varepsilon_{i,t}|\leq|v_{i}(t)|+\varepsilon\leq|v_{i}(t)|+\lambda\leq 2|v_{i}(t)|$ (3)
Also, for the auxiliary item $(m+1)$, the assigned value is sufficiently small: $\widehat{v}_{i}(m+1)=\lambda/2$ for each $i\in[n]$. This observation and equation (3) gives us the desired bound $\max_{{i\in[n],\ t\in[m]}}\ |v_{i}(t)|\ \geq\ \frac{1}{2}\max_{{i\in[n],\ t\in[\widehat{m}]}}\ |\widehat{v}_{i}(t)|$. ∎
The following two lemmas provide relevant properties of the constructed instance $\widehat{\mathcal{I}}$. Specifically, Lemma 2 shows that all the items in $\widehat{\mathcal{I}}$ can be categorized into three types: goods, chores, or non-negatively valued. Then, Lemma 3 establishes that valuations $\widehat{v}_{i}(\cdot)$—obtained by perturbing the given values—satisfy a useful non-degeneracy property.
Lemma 2.
[…]
Proof.
[…]
We will show that the following containments hold between $G^{\prime},C^{\prime},G_{0}^{\prime}$ and the subsets $G,C,G_{0}$ (as defined in the lemma): $G=G^{\prime}\cup\{(m+1)\}$ along with $C=C^{\prime}$ and $G_{0}=G^{\prime}_{0}$. Since $\widehat{m}]=[m]\cup\{(m+1)\}$, these containments establish the lemma.
The sign preservations stated in Proposition 1 lead to the stated containments: For each item $t\in G^{\prime}$, we have $v_{i}(t)>0$ for all the agents $i\in[n]$. Therefore, $\widehat{v}_{i}(t)>0$ for all agents $i$, i.e., as desired, $t\in G$. Also, for each item $t^{\prime}\in C^{\prime}$, it holds that $v_{j}(t^{\prime})<0$, for all the agents $j\in[n]$. This implies $\widehat{v}_{j}(t^{\prime})<0$ for each $j\in[n]$ and, hence, $t^{\prime}\in C$. Similarly, using sign preservation, we obtain that each non-negatively valued item in $\mathcal{I}$ continues to be non-negatively valued in $\widehat{\mathcal{I}}$; $s\in G^{\prime}_{0}$ implies $s\in G_{0}$.
Finally, note that for item $(m+1)$, we have $\widehat{v}_{i}(m+1)=\lambda/2>0$, for each $i\in[n]$, hence, the stated containment, $(m+1)\in G$, holds.
Therefore, the set of items $\widehat{m}$] partitions into goods $G$, chores $C$, and non-negatively valued items $G_{0}$. The lemma stands proved. ∎
Lemma 3.
In the instance $\widehat{\mathcal{I}}=\langle[n],[\widehat{m}],\{\widehat{v}_{i}\}_{i=1}^{n}\rangle$ and for any simple cycle $C=(i_{1},t_{1},i_{2},t_{2},\ldots,i_{k},t_{k},i_{1})$ in the complete bipartite graph $K_{n,\widehat{m}}$, [^6] Note that, since the cycles $C$ are simple, any agent $i_{\ell}\in[n]$ and any item $t_{\ell}\in[\widehat{m}$] can appear at most once in $C$. with $\widehat{v}_{i_{\ell}}(t_{\ell}),\widehat{v}_{i_{\ell+1}}(t_{\ell})\neq 0$, for each index $\ell\in[k]$, it holds that $\prod_{\ell=1}^{k}\left(\frac{\widehat{v}_{i_{\ell+1}}(t_{\ell})}{\widehat{v}_{i_{\ell}}(t_{\ell})}\right)\neq 1$; here, the indices are cyclically mapped, i.e., $i_{k+1}=i_{1}$.
Proof.
Fix any simple cycle $C=(i_{1},t_{1},i_{2},t_{2},\ldots,i_{k},t_{k},i_{1})$ in $K_{n,\widehat{m}}$ with $\widehat{v}_{i_{\ell}}(t_{\ell}),\widehat{v}_{i_{\ell+1}}(t_{\ell})\neq 0$, for each $\ell\in[k]$. Since cycle $C$ contains at least two items and item $(m+1)$ can appear at most once in the simple cycle, there exists an item $t_{x}$ in $C$ such that $t_{x}\neq(m+1)$. For such an item $t_{x}$, the values $\widehat{v}_{i_{x}}(t_{x})$ and $\widehat{v}_{i_{x+1}}(t_{x})$ are nonzero and satisfy $\widehat{v}_{i_{x}}(t_{x})=v_{i_{x}}(t_{x})-\varepsilon_{i_{x},t_{x}}$ along with $\widehat{v}_{i_{x+1}}(t_{x})=v_{i_{x+1}}(t_{x})-\varepsilon_{i_{x+1},t_{x}}$. Here, $\varepsilon_{i_{x},t_{x}}$ and $\varepsilon_{i_{x+1},t_{x}}$ are continuous random variables independently drawn from the uniform distribution over $0,\varepsilon]$. Hence, the product $\prod_{\ell=1}^{k}\left(\frac{\widehat{v}_{i_{\ell+1}}(t_{\ell})}{\widehat{v}_{i_{\ell}}(t_{\ell})}\right)$ is obtained by multiplying independent [^7] Recall that $\varepsilon_{i,t}$s are drawn independently. continuous random variables. Therefore, the product itself is a continuous random variable. That is, $\prod_{\ell=1}^{k}\left(\frac{\widehat{v}_{i_{\ell+1}}(t_{\ell})}{\widehat{v}_{i_{\ell}}(t_{\ell})}\right)\neq 1$ with probability one.
[…]
4.2 Reduction to Non-degenerate Instance $\widehat{\mathcal{I}}$
We will show that, to prove Theorem 1 it suffices to focus on the non-degenerate instance $\widehat{\mathcal{I}}$. Towards this, we first set parameter $\eta>0$ as
$\eta\coloneqq\frac{\lambda}{2mn\max\limits_{i\in[n],\ t\in[\widehat{m}]}|\widehat{v}_{i}(t)|}.$ (4)
The definition of $\eta$ leads to following relation between this parameter and $\varepsilon$:
$\frac{\eta\omega}{4m}$ $=\frac{\omega}{4m}\left(\frac{\lambda}{2mn\max\limits_{i\in[n],\ t\in[\widehat{m}]}|\widehat{v}_{i}(t)|}\right)$
$\geq\frac{\omega}{4m}\left(\frac{\lambda}{4mn\max\limits_{i\in[n],\ t\in[m]}|v_{i}(t)|}\right)$ (via Proposition 1)
$=\frac{\lambda\omega}{16m^{2}n\max\limits_{i\in[n],\ t\in[m]}|v_{i}(t)|}$
$>\varepsilon$ (5)
Here, the last (strict) inequality follows from the definition of $\varepsilon$ (Equation 1).
For instance $\widehat{\mathcal{I}}$ and constant $\eta>0$, we define the following primal and dual linear programs parameterized by weight vectors $w\in\Delta_{n-1}$. In particular, the linear program $\textbf{{LP}}(w,\eta)$ maximizes the weighted social welfare across all fractional allocations.
[…]
Note that, for any weight vector $w\in\Delta_{n-1}$, the linear program $\textbf{{LP}}(w,\eta)$ always admits an integral optimal solution. Such an optimum corresponds to an integral allocation $\mathcal{A}=(A_{1},\ldots,A_{n})\in\Pi_{n}([m])$ that can be obtained by assigning each item $t$ to an agent $i\in\operatorname*{arg\,max}_{j\in[n]}\ (w_{j}+\eta)\widehat{v}_{j}(t)$. We will write $\mathcal{A}\in\mathrm{OPT}(\textbf{{LP}}(w,\eta))$ to denote the weighted welfare optimality of $\mathcal{A}$. In fact, there can be multiple (integral) allocations in the set of optimal solutions $\mathrm{OPT}(\textbf{LP}(w,\eta))$. By contrast, for each $w\in\Delta_{n-1}$, the dual linear program $\textbf{{Dual-LP}}(w,\eta)$ has a unique optimal solution given by $p_{t}=\max_{i\in[n]}\ \bigl(w_{i}+\eta\bigr)\widehat{v}_{i}(t)$; we will denote this fact by writing $p=\mathrm{OPT}(\textbf{{Dual-LP}}(w,\eta))$ and refer to $p\in\mathbb{R}^{\widehat{m}}$ as the vector of optimal prices.
The proposition below shows that the optimal prices are always nonzero.
Proposition 2.
For any weight vector $w\in\Delta_{n-1}$, if $p=\mathrm{OPT}(\textbf{{Dual-LP}}(w,\eta))$, then $p_{t}>0$ for each item $t\in G\cup G_{0}$ and $p_{t}<0$ for each $t\in C$.
Proof.
The optimal prices $p=\mathrm{OPT}(\textbf{{Dual-LP}}(w,\eta))$ satisfy $p_{t}=\max_{i\in[n]}\ \bigl(w_{i}+\eta\bigr)\widehat{v}_{i}(t)$ for each item $t\in[\widehat{m}$]. Since $\eta>0$ and $w\in\Delta_{n-1}$, it holds that the weight $\bigl(w_{i}+\eta\bigr)>0$ for each $i\in[n]$.
Now, note that, for each item $t\in G\cup G_{0}$ (by definition of the item subsets $G$ and $G_{0}$), we have $\max_{i\in[n]}\ \widehat{v}_{i}(t)>0$. Hence, for each $t\in G\cup G_{0}$, the optimal price $p_{t}=\max_{i\in[n]}\ \bigl(w_{i}+\eta\bigr)\widehat{v}_{i}(t)>0$. On the other hand, for each item $s\in C$, the value $\widehat{v}_{i}(s)<0$ for every agent $i\in[n]$. Hence, $p_{s}=\max_{i\in[n]}\bigl(w_{i}+\eta\bigr)\widehat{v}_{i}(s)<0$ for items $s\in C$. This completes the proof of the proposition. ∎
The following two lemmas establish that if an allocation $\mathcal{A}$ is fair and welfare-maximizing in the non-degenerate instance $\widehat{\mathcal{I}}$, then $\mathcal{A}$ (with the auxiliary item $(m+1)$ removed) is fair and $\mathrm{PO}$ in the given instance $\mathcal{I}$. That is, these two lemmas ensure that the existence of a fair and welfare-maximizing allocation in $\widehat{\mathcal{I}}$ implies our main result (Theorem 1) for $\mathcal{I}$.
Lemma 4.
If an allocation $\mathcal{A}=(A_{1},\ldots,A_{n})$ is $\mathrm{IEF}1$ in the non-degenerate instance $\widehat{\mathcal{I}}=\langle[n],[\widehat{m}],\{\widehat{v}_{i}\}_{i=1}^{n}\rangle$, then the allocation $(A_{1}\setminus\{(m+1)\},A_{2}\setminus\{(m+1)\},\ldots,A_{n}\setminus\{(m+1)\})$ is $\mathrm{IEF}1$ in the given instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$.
Lemma 5.
If an allocation $\mathcal{A}=(A_{1},\ldots,A_{n})\in\mathrm{OPT}(\textbf{{LP}}(w,\eta))$, for any $w\in\Delta_{n-1}$, then the allocation $(A_{1}\setminus\{(m+1)\},A_{2}\setminus\{(m+1)\},\ldots,A_{n}\setminus\{(m+1)\})$ is $\mathrm{PO}$ in the instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$.
Upon proving these lemmas, the subsequent goal of the paper will be to establish the existence of a $w^{*}\in\Delta_{n-1}$ and an allocation $\mathcal{A}^{*}\in\mathrm{OPT}(\textbf{LP}(w^{*},\eta))$ that is $\mathrm{IEF}1$ in the non-degenerate instance $\widehat{\mathcal{I}}$. As mentioned previously, Lemma 4 and 5 will then imply that $(A^{*}_{1}\setminus\{(m+1)\},A^{*}_{2}\setminus\{(m+1)\},\ldots,A^{*}_{n}\setminus\{(m+1)\})$ is $\mathrm{IEF}1$ and $\mathrm{PO}$ for the given instance $\mathcal{I}$, thereby proving Theorem 1.
Proof of Lemma 4.
Allocation $\mathcal{A}=(A_{1},\ldots,A_{n})$ is $\mathrm{IEF}1$ in instance $\widehat{\mathcal{I}}$ and, hence, for each agent $i\in[n]$ there exists a subset $S_{i}\subseteq[m]$, with $|S_{i}|\leq 1$ and the property that $\widehat{v}_{i}(A_{i}\triangle S_{i})\geq\max_{j\in[n]}\ \widehat{v}_{i}(A_{j})$. We will show that the inequality $\widehat{v}_{i}(A_{i}\triangle S_{i})\geq\max_{j\in[n]}\widehat{v}_{i}(A_{j})$ implies the following bound (under the given valuation $v_{i}$s)
$v_{i}\Big((A_{i}\setminus\{m+1\})\triangle(S_{i}\setminus\{m+1\})\Big)\allowbreak\geq\max_{j\in[n]}v_{i}\big(A_{j}\setminus\{m+1\}\big).$ (6)
Since $|S_{i}\setminus\{m+1\}|\leq 1$, this would imply that the allocation $(A_{1}\setminus\{m+1\},A_{2}\setminus\{m+1\},\ldots,A_{n}\setminus\{m+1\})$ is $\mathrm{IEF}1$ for the instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$, thereby establishing the lemma. To prove equation (6), we start with inequality $\widehat{v}_{i}(A_{i}\triangle S_{i})\geq\max_{j\in[n]}\widehat{v}_{i}(A_{j})$, and, for each $j\in[n]$, obtain:
$0$ $\leq\widehat{v}_{i}(A_{i}\triangle S_{i})-\widehat{v}_{i}(A_{j})$
$\leq\widehat{v}_{i}\big((A_{i}\triangle S_{i})\setminus\{m+1\}\big)-\widehat{v}_{i}(A_{j}\setminus\{m+1\})+\widehat{v}_{i}(m+1)$ (given that $\widehat{v}_{i}(m+1)>0$)
$\leq v_{i}\big((A_{i}\triangle S_{i})\setminus\{m+1\}\big)-v_{i}(A_{j}\setminus\{m+1\})+\varepsilon|A_{j}\setminus\{m+1\}|+\widehat{v}_{i}(m+1)$ (since $v_{i}(t)-\varepsilon\leq\widehat{v}_{i}(t)\leq v_{i}(t)$, for each $t$)
$\leq v_{i}\big((A_{i}\triangle S_{i})\setminus\{m+1\}\big)-v_{i}(A_{j}\setminus\{m+1\})+\varepsilon m+\frac{\lambda}{2}$ ($\widehat{v}_{i}(m+1)=\lambda/2$)
$<v_{i}\big((A_{i}\triangle S_{i})\setminus\{m+1\}\big)-v_{i}(A_{j}\setminus\{m+1\})+\frac{\lambda}{8m}\ m+\frac{\lambda}{2}$ ($\varepsilon<\frac{\lambda}{8m}$; Equation 2)
$\leq v_{i}\Big((A_{i}\setminus\{m+1\})\triangle(S_{i}\setminus\{m+1\})\Big)-v_{i}(A_{j}\setminus\{m+1\})+\lambda$
Hence, $v_{i}(A_{j}\setminus\{m+1\})-v_{i}\Big((A_{i}\setminus\{m+1\})\triangle(S_{i}\setminus\{m+1\})\Big)<\lambda$. This strict inequality implies that, in fact, $v_{i}(A_{j}\setminus\{m+1\})-v_{i}\Big((A_{i}\setminus\{m+1\})\triangle(S_{i}\setminus\{m+1\})\Big)\leq 0$; otherwise, the bound $0<v_{i}(A_{j}\setminus\{m+1\})-v_{i}\Big((A_{i}\setminus\{m+1\})\triangle(S_{i}\setminus\{m+1\})\Big)<\lambda$ would contradict the definition of $\lambda$. Using $v_{i}(A_{j}\setminus\{m+1\})-v_{i}\Big((A_{i}\setminus\{m+1\})\triangle(S_{i}\setminus\{m+1\})\Big)\leq 0$ for all agents $j\in[n]$, we obtain equation (6). This completes the proof of the lemma. ∎
Next, we establish Lemma 5.
Proof of Lemma 5.
We begin by defining an intermediate instance $\mathcal{I}^{\prime}=\langle[n],[m+1],\{v^{\prime}_{i}\}_{i=1}^{n}\rangle$ wherein $v^{\prime}_{i}(t)=v_{i}(t)$, for all $i\in[n]$ and $t\in[m]$, and $v^{\prime}_{i}(m+1)=\lambda/2$ for each $i\in[n]$. That is, we construct instance $\mathcal{I}^{\prime}$ by adding item $(m+1)$ to $\mathcal{I}$ and keeping the valuations of the given $m$ items unchanged. Instance $\mathcal{I}^{\prime}$ is defined for the purposes of analysis: We will first show that $\mathcal{A}$ is $\mathrm{PO}$ in $\mathcal{I}^{\prime}$. Then, we will conclude the proof by showing that $(A_{1}\setminus\{m+1\},A_{2}\setminus\{m+1\},\ldots,A_{n}\setminus\{m+1\})$ is $\mathrm{PO}$ in $\mathcal{I}$.
Assume, towards a contradiction, that there exists an allocation $\mathcal{B}=(B_{1},\ldots,B_{n})$ that Pareto dominates $\mathcal{A}$ in $\mathcal{I}^{\prime}$. That is, $v^{\prime}_{i}(B_{i})\geq v^{\prime}_{i}(A_{i})$, for all $i\in[n]$, and $v^{\prime}_{a}(B_{a})>v^{\prime}_{a}(A_{a})$ for some $a\in[n]$. These inequalities imply that
$0$ $<\sum_{i=1}^{n}v^{\prime}_{i}(B_{i})-\sum_{i=1}^{n}v^{\prime}_{i}(A_{i})$
$=\frac{\lambda}{2}+\sum_{i=1}^{n}v^{\prime}_{i}(B_{i}\setminus\{m+1\})-\frac{\lambda}{2}-\sum_{i=1}^{n}v^{\prime}_{i}(A_{i}\setminus\{m+1\})$ ($v^{\prime}_{j}(m+1)=\lambda/2$ for all $j$)
$=\sum_{i=1}^{n}v_{i}(B_{i}\setminus\{m+1\})-\sum_{i=1}^{n}v_{i}(A_{i}\setminus\{m+1\})$ (7)
That is, $\sum_{i=1}^{n}v_{i}(B_{i}\setminus\{m+1\})-\sum_{i=1}^{n}v_{i}(A_{i}\setminus\{m+1\})>0$. Using this strict inequality and the definition of $\omega$, we obtain
$\omega\leq\sum_{i=1}^{n}v_{i}(B_{i}\setminus\{m+1\})-\sum_{i=1}^{n}v_{i}(A_{i}\setminus\{m+1\})=\sum_{i=1}^{n}v^{\prime}_{i}(B_{i})-\sum_{i=1}^{n}v^{\prime}_{i}(A_{i})$
For the last equality, we use the fact that $v^{\prime}_{j}(m+1)=\lambda/2$ for all $j$.
Next, we use the welfare optimality of $\mathcal{A}$ in $\widehat{\mathcal{I}}$ (and under the weight vector $w$) to arrive at a contradiction. In particular, $\mathcal{A}\in\mathrm{OPT}(\textbf{{LP}}(w,\eta))$ implies that $\sum_{i=1}^{n}(w_{i}+\eta)\widehat{v}_{i}(B_{i})\leq\sum_{i=1}^{n}(w_{i}+\eta)\widehat{v}_{i}(A_{i})$. For each subset $S\subseteq[\widehat{m}$] and each agent $i$, it holds that $v^{\prime}_{i}(S)-\varepsilon|S|\leq\widehat{v}_{i}(S)\leq v^{\prime}_{i}(S)$. Therefore, we have $\sum_{i=1}^{n}(w_{i}+\eta)(v^{\prime}_{i}(B_{i})-\varepsilon|B_{i}|)\leq\sum_{i=1}^{n}(w_{i}+\eta)v^{\prime}_{i}(A_{i})$. This inequality reduces to
$\sum_{i=1}^{n}(w_{i}+\eta)(v^{\prime}_{i}(B_{i})-v^{\prime}_{i}(A_{i}))$ $\leq\varepsilon\sum_{i=1}^{n}(w_{i}+\eta)|B_{i}|$
$=\varepsilon\sum_{i=1}^{n}w_{i}|B_{i}|\ +\ \varepsilon\eta\sum_{i=1}^{n}|B_{i}|$
$\leq\varepsilon(m+1)\sum_{i=1}^{n}w_{i}\ +\ \varepsilon\eta\sum_{i=1}^{n}|B_{i}|$ (since $w_{i}\geq 0$ and $|B_{i}|\leq m+1$)
$=\varepsilon(m+1)+\varepsilon\eta(m+1)$ (since $\sum_{i}w_{i}=1$ and $\sum_{i}|B_{i}|=m+1$)
$=\varepsilon(m+1)(1+\eta)$
$\leq 2\varepsilon(m+1)$ (since $1+\eta<2$)
$\leq 4\varepsilon m$ (9)
We can lower bound the left-hand-side of inequality (9) as follows
$\sum_{i=1}^{n}(w_{i}+\eta)(v^{\prime}_{i}(B_{i})-v^{\prime}_{i}(A_{i}))$ $\geq\eta\sum_{i=1}^{n}(v^{\prime}_{i}(B_{i})-v^{\prime}_{i}(A_{i}))$ (since $v^{\prime}_{i}(B_{i})\geq v^{\prime}_{i}(A_{i})$ and $w_{i}\geq 0$, for each $i$)
$\geq\eta\omega$ (via (4.2))
The last inequality and equation (9) imply $\eta\omega\leq 4\varepsilon m$, i.e., $\frac{\eta\omega}{4m}\leq\varepsilon$. This bound, however, contradicts the strict inequality (5). Therefore, by way of contradiction, we obtain that no allocation $\mathcal{B}$ Pareto dominates $\mathcal{A}\in\mathrm{OPT}(\textbf{{LP}}(w,\eta))$ in the instance $\mathcal{I}^{\prime}$, i.e., $\mathcal{A}$ is $\mathrm{PO}$ in $\mathcal{I}^{\prime}$.
Finally, we will use the Pareto optimality of $\mathcal{A}$ in $\mathcal{I}^{\prime}$ to show that the allocation $(A_{1}\setminus\{m+1\},\ldots,A_{n}\setminus\{m+1\})$ is $\mathrm{PO}$ in the given instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$. Assume, towards a contradiction, that allocation $(D_{1},\ldots,D_{n})$ Pareto dominates $(A_{1}\setminus\{m+1\},\ldots,A_{n}\setminus\{m+1\})$ in $\mathcal{I}$. Write $a\in[n]$ to denote the agent that receives the auxiliary item $(m+1)$ in allocation $\mathcal{A}$, i.e., $m+1\in A_{a}$. Note that the allocation $(A_{1}\setminus\{m+1\},\ldots,A_{a}\setminus\{m+1\},\ldots,A_{n}\setminus\{m+1\})$ is same as $(A_{1},\ldots,A_{a}\setminus\{m+1\},\ldots,A_{n})$. The fact that $(D_{1},\ldots,D_{n})$ Pareto dominates $(A_{1},\ldots,A_{a}\setminus\{m+1\},\ldots,A_{n})$ implies that $(D_{1},\ldots,D_{a}\cup\{m+1\},\ldots,D_{n})$ Pareto dominates $\mathcal{A}$ in instance $\mathcal{I}^{\prime}$. Since this contradicts the already established Pareto optimality of $\mathcal{A}$ in $\mathcal{I}^{\prime}$, we obtain that $(A_{1}\setminus\{m+1\},\ldots,A_{n}\setminus\{m+1\})$ is $\mathrm{PO}$ in the instance $\mathcal{I}$. The lemma stands proved. ∎
In light of Lemmas 4 and 5, our goal (in Sections 5 and 6) will be to prove the following theorem.
Theorem 6.
There exists a vector $w^{*}\in\Delta_{n-1}$ and an associated allocation $\mathcal{A}^{*}\in\mathrm{OPT}(\textbf{{LP}}(w^{*},\eta))$ with the property that $\mathcal{A}^{*}$ is $\mathrm{IEF}1$ in the constructed instance $\widehat{\mathcal{I}}$.
As stated previously, our main result, Theorem 1, follows directly from Theorem 6. For completeness, we provide a proof here.
Proof of Theorem 1.
Let $\mathcal{A}^{*}=(A^{*}_{1},\ldots,A^{*}_{n})$ be the allocation whose existence is guaranteed by Theorem 6. By Lemmas 4 and 5, the allocation $(A^{*}_{1}\setminus\{m+1\},A^{*}_{2}\setminus\{m+1\},\ldots,A^{*}_{n}\setminus\{m+1\})$ is $\mathrm{IEF}1$ and $\mathrm{PO}$ in the given instance $\mathcal{I}$. Therefore, we obtain the stated fairness and efficiency guarantee for indivisible mixed manna. ∎
5 Applying the KKM Theorem
To prove Theorem 6, we will invoke the Knaster-Kuratowski-Mazurkiewicz (KKM) Theorem. The theorem asserts that any $n$ closed subsets of the (n1)(n-1)-dimensional simplex $\Delta_{n-1}$ that satisfy a covering condition necessarily have a nonempty intersection.
Theorem 7.
[knaster1929beweis] Let $C_{1},C_{2},\ldots,C_{n}\subseteq\Delta_{n-1}$ be closed subsets of $\Delta_{n-1}$ with the property that for each vector $w\in\Delta_{n-1}$ there is an index $i\in\mathrm{supp}(w)$ satisfying $w\in C_{i}$. Then, there exists a vector $w^{*}\in\cap_{i=1}^{n}C_{i}$.
To apply Theorem 7, we will appropriately define the subsets $C_{i}\subseteq\Delta_{n-1}$ and show that the vector $w^{*}$ guaranteed in the KKM Theorem satisfies Theorem 6.
We define $C_{i}$s considering the previously-mentioned primal and dual linear programs, $\textbf{{LP}}(w,\eta)$ and $\textbf{{Dual-LP}}(w,\eta)$, which are parameterized by weight vectors $w\in\Delta_{n-1}$. Recall that $p=\mathrm{OPT}(\textbf{{Dual-LP}}(w,\eta))$ denotes the optimal price vector and $\mathrm{OPT}(\textbf{{LP}}(w,\eta))$ contains the weighted welfare-maximizing (integral) allocations. For any subset of items $S\subseteq[\widehat{m}$], we will write $p(S)$ to denote the total price of the items in $S$, i.e., $p(S)=\sum_{t\in S}p_{t}$.
A key construct here is price envy-freeness. Specifically, we say that an agent $i$ is price envy-free in an allocation $\mathcal{B}=(B_{1},\ldots,B_{n})$ and under prices $p^{\prime}\in\mathbb{R}^{\widehat{m}}$ if $i$’s bundle has a price at least as much as that of any other agent: $p^{\prime}(B_{i})\geq p^{\prime}(B_{j})$ for all $j\in[n]$. For each agent $i\in[n]$, we define $C_{i}$ to be the set of weights $w\in\Delta_{n-1}$ for which there is an optimal allocation $\mathcal{A}\in\mathrm{OPT}(\textbf{{LP}}(w,\eta))$ wherein $i$ is price envy-free, under the optimal prices $p=\mathrm{OPT}(\textbf{{Dual-LP}}(w,\eta))$. Formally,
[…]
The following two lemmas show that the sets $C_{1},\ldots,C_{n}$ (defined in equation (10)) satisfy the conditions of the KKM Theorem.
Lemma 8 (Closedness).
For each $i\in[n]$, the set $C_{i}\subseteq\Delta_{n-1}$ is closed.
Proof.
Fix any agent $i\in[n]$. Consider any convergent sequence $\{w^{k}\}_{k\in\mathbb{N}}$ in $C_{i}$ and $\overline{w}\in\Delta_{n-1}$ be its limit point. We will show that $\overline{w}\in C_{i}$ and, hence, obtain that the $C_{i}$ is closed.
Since the total number of allocations are finite, there must be an infinite subsequence $\{\widetilde{w}^{k}\}_{k\in\mathbb{N}}$ and a particular allocation $\widetilde{\mathcal{A}}=(\widetilde{A}_{1},\ldots,\widetilde{A}_{n})$ such that, for each $k\in\mathbb{N}$, allocation $\widetilde{\mathcal{A}}\in\mathrm{OPT}(\textbf{{LP}}(\widetilde{w}^{k},\eta))$ and $i$ is price envy-free here with optimal prices $\widetilde{p}^{k}=\mathrm{OPT}(\textbf{{Dual-LP}}(\widetilde{w}^{k},\eta))$.
We first note that $\widetilde{\mathcal{A}}$ is a welfare-maximizing allocation even for the limit point $\overline{w}$.
Claim 1.
$\widetilde{\mathcal{A}}\in\mathrm{OPT}(\textbf{{LP}}(\overline{w},\eta))$.
Proof.
Write $\mathrm{SW}(\mathcal{B},w)$ to denote the weighted social welfare of allocation $\mathcal{B}=(B_{1},\ldots,B_{n})$, [^8] Recall that $\eta$ is a fixed parameter. i.e., $\mathrm{SW}(\mathcal{B},w)\coloneqq\sum_{i=1}^{n}(w_{i}+\eta)\widehat{v}_{i}(B_{i})$. In addition, write $\mathrm{SW}^{*}(w)$ to denote that optimal value of the linear program $\textbf{{LP}}(w,\eta)$. Note that the optimal value satisfies $\mathrm{SW}^{*}(w)=\max_{\mathcal{B}\in\Pi_{n}([m])}\ \mathrm{SW}(\mathcal{B},w)$.
Furthermore, for any fixed allocation $\mathcal{B}\in\Pi_{n}([m])$, the function $\mathrm{SW}(\mathcal{B},w)$ is continuous in $w$. Hence, $\mathrm{SW}^{*}(w)$—a maximum of continuous functions—is also continuous. Now, using the sequential criterion of continuity, we obtain that $\mathrm{SW}(\widetilde{\mathcal{A}},\widetilde{w}^{k})\to\mathrm{SW}(\widetilde{\mathcal{A}},\overline{w})$ as $k\to\infty$. In addition, $\mathrm{SW}^{*}(\widetilde{w}^{k})\to\mathrm{SW}^{*}(\overline{w})$ as $k\to\infty$. The two sequences here, $\{\mathrm{SW}(\widetilde{\mathcal{A}},\widetilde{w}^{k})\}_{k}$ and $\{\mathrm{SW}^{*}(\widetilde{w}^{k})\}_{k}$, are the same, since $\widetilde{\mathcal{A}}\in\mathrm{OPT}(\textbf{{LP}}(\widetilde{w}^{k},\eta))$ for each $k$. Therefore, $\mathrm{SW}(\widetilde{\mathcal{A}},\overline{w})=\mathrm{SW}^{*}(\overline{w})$, i.e., allocation $\widetilde{\mathcal{A}}$ is welfare-maximizing at $\overline{w}$ as well: $\widetilde{\mathcal{A}}\in\mathrm{OPT}(\textbf{{LP}}(\overline{w},\eta))$. This completes the proof of the claim. ∎
Write $\overline{p}=\mathrm{OPT}(\textbf{{Dual-LP}}(\overline{w},\eta))$ to denote the optimal price vector at the limit point $\overline{w}$. The claim below shows that the allocation $\widetilde{\mathcal{A}}=(\widetilde{A}_{1},\ldots,\widetilde{A}_{n})$ ensures price envy-freeness for the agent $i$ under $\overline{p}$.
Claim 2.
Agent $i\in\operatorname*{arg\,max}_{j\in[n]}\ \overline{p}(\widetilde{A}_{j})$.
Proof.
Recall that for any weight vector $w^{\prime}$, the optimal prices $p^{\prime}=\mathrm{OPT}(\textbf{{Dual-LP}}(w,\eta))$ satisfy $p^{\prime}_{t}=\max_{i\in[n]}\ (w^{\prime}_{i}+\eta)\widehat{v}_{i}(t)$ for each item $t\in[\widehat{m}$]. That is, for each item $t$, the optimal price is a continuous function of the weight. Hence, using the sequential criterion of continuity again, we obtain that the optimal price vectors $\widetilde{p}^{k}=\mathrm{OPT}(\textbf{{Dual-LP}}(\widetilde{w}^{k},\eta))$ converge (component-wise) to $\overline{p}$. Furthermore, for each bundle $\widetilde{A}_{j}$ in the allocation $\widetilde{\mathcal{A}}=(\widetilde{A}_{1},\ldots,\widetilde{A}_{n})$, we have $\widetilde{p}^{k}\left(\widetilde{A}_{j}\right)\to\overline{p}\left(\widetilde{A}_{j}\right)$ as $k\to\infty$.
Next, note that, for each $k\in\mathbb{N}$, since agent $i$ is price-envy free in $\widetilde{\mathcal{A}}$ (under $\widetilde{p}^{k}$), we have $\widetilde{p}^{k}\left(\widetilde{A}_{i}\right)\geq\widetilde{p}^{k}\left(\widetilde{A}_{j}\right)$ for every $j\in[n]$. Therefore, the order-preserving property of limits gives us $\overline{p}\left(\widetilde{A}_{i}\right)\geq\overline{p}\left(\widetilde{A}_{j}\right)$ for each $j$. The claim stands proved. ∎
Claims 1 and 2 imply that, for the limit point $\overline{w}$, the allocation $\widetilde{\mathcal{A}}\in\mathrm{OPT}(\textbf{{LP}}(\overline{w},\eta))$ yields price envy-freeness for agent $i$ under $\overline{p}$. Therefore, the limit point $\overline{w}\in C_{i}$ (see equation (10)) and, as desired, we obtain that $C_{i}$ is closed. ∎
Lemma 9 (KKM Covering).
For each $w\in\Delta_{n-1}$ there exists an $i\in\mathrm{supp}(w)$ such that $w\in C_{i}$.
Proof.
Fix any weight vector $w\in\Delta_{n-1}$. We consider two cases, based on whether $\mathrm{supp}(w)=[n]$ or $\mathrm{supp}(w)\subsetneq[n]$.
Case I: $\mathrm{supp}(w)=[n]$. Consider any allocation $\mathcal{A}\in\mathrm{OPT}(\textbf{{LP}}(w,\eta))$ and write $p=\mathrm{OPT}(\textbf{{Dual-LP}}(w,\eta))$. Also, write $i^{\prime}\in\operatorname*{arg\,max}_{j\in[n]}p(A_{j})$. In the current case $\mathrm{supp}(w)=[n]$ and, hence, $i^{\prime}\in\mathrm{supp}(w)$. Additionally, we have $w\in C_{i^{\prime}}$, since $i^{\prime}\in\operatorname*{arg\,max}_{j\in[n]}p(A_{j})$ (see equation (10)). Hence, for vectors $w$ with full support, the KKM covering condition holds.
Case II: $\mathrm{supp}(w)\subsetneq[n]$. In this case $n]\setminus\mathrm{supp}(w)\neq\emptyset$. Write $p=\mathrm{OPT}(\textbf{{Dual-LP}}(w,\eta))$ and consider any allocation $\mathcal{A}\in\mathrm{OPT}(\textbf{{LP}}(w,\eta))$. We will show that there exists an $i^{\prime}\in\mathrm{supp}(w)$ with the property that $i^{\prime}\in\operatorname*{arg\,max}_{j\in[n]}\ p(A_{j})$, i.e., $i^{\prime}$ (with $w_{i^{\prime}}>0$) is price envy-free. For such an agent $i^{\prime}$, we have $w\in C_{i^{\prime}}$, and, hence, we will obtain KKM covering in the current case as well.
Write $H\coloneqq\{i\in[n]\mid w_{i}\geq 1/n\}$. This set $H\neq\emptyset$, since $\sum_{i=1}^{n}w_{i}=1$. Also, the set of agents $[n]$ partitions into three disjoint subsets: $n]\setminus\mathrm{supp}(w)$, $H$, and $\mathrm{supp}(w)\setminus H$.
Furthermore, note that in the welfare-maximizing allocation $\mathcal{A}\in\mathrm{OPT}(\textbf{{LP}}(w,\eta))$, each item $t\in[\widehat{m}$] is assigned to an agent in $\operatorname*{arg\,max}_{i\in[n]}\ (w_{i}+\eta)\widehat{v}_{i}(t)$. In particular, the auxiliary item $(m+1)$ must be assigned (in $\mathcal{A}$) to an agent in $\operatorname*{arg\,max}_{i\in[n]}w_{i}$; recall that, by construction, item $(m+1)$ has the same value $\widehat{v}_{i}(m+1)=\lambda/2>0$ for all the agents $i$. Write $h\in\operatorname*{arg\,max}_{i\in[n]}w_{i}$ to denote the agent that receives $(m+1)$ in $\mathcal{A}$. That is, $h$ has a highest-valued component in $w$ and item $(m+1)\in A_{h}$. Also, we have $h\in H\subseteq\mathrm{supp}(w)$.
To establish the existence of price envy-free agent $i^{\prime}\in\mathrm{supp}(w)$, we will show that $\max\limits_{j\in[n]\setminus\mathrm{supp}(w)}p(A_{j})\leq p(A_{h})$. Since $h\in\mathrm{supp}(w)$, the last inequality extends to
$\max\limits_{j\in[n]\setminus\mathrm{supp}(w)}p(A_{j})\leq p(A_{h})\leq\max\limits_{i\in\mathrm{supp}(w)}p(A_{i})$ (11)
Note that inequality (11) implies that there exists an agent $i^{\prime}\in\mathrm{supp}(w)$ with a maximum-price bundle, $i^{\prime}\in\operatorname*{arg\,max}_{i\in[n]}p(A_{i})$, i.e., as desired, there exists a price envy-free agent $i^{\prime}\in\mathrm{supp}(w)$.
Hence, to complete the analysis of the current case, it remains to prove that $\max\limits_{j\in[n]\setminus\mathrm{supp}(w)}p(A_{j})\leq p(A_{h})$. Towards this, we will establish the following two inequalities
$\max\limits_{j\in[n]\setminus\mathrm{supp}(w)}p(A_{j})$ $\leq\frac{\lambda}{2n}$ (12)
$\frac{\lambda}{2n}$ $\leq p(A_{h})$ (13)
Indeed, inequalities (12) and (13) imply the stated bound: $\max\limits_{j\in[n]\setminus\mathrm{supp}(w)}p(A_{j})\leq\frac{\lambda}{2n}\leq p(A_{h})$.
Establishing Inequality (12). As mentioned above, in the welfare-maximizing allocation $\mathcal{A}\in\mathrm{OPT}(\textbf{{LP}}(w,\eta))$, each item $t\in[\widehat{m}$] is assigned to an agent in $\operatorname*{arg\,max}_{i\in[n]}\ (w_{i}+\eta)\widehat{v}_{i}(t)$. In particular, for each agent $b\in[n]\setminus\mathrm{supp}(w)$ and each item $s\in A_{b}$, it holds that $p_{s}=\max_{i\in[n]}(w_{i}+\eta)\widehat{v}_{i}(t)=\eta\ \widehat{v}_{b}(s)$; recall that $w_{b}=0$, since $b\in[n]\setminus\mathrm{supp}(w)$. Since the auxiliary item $(m+1)$ is assigned to the agent $h\notin[n]\setminus\mathrm{supp}(w)$, for each $b\in[n]\setminus\mathrm{supp}(w)$, we have $|A_{b}|\leq m$. These observations give us
$p(A_{b})=\sum_{s\in A_{b}}p_{s}=\sum_{s\in A_{b}}\eta\ \widehat{v}_{b}(s)\leq|A_{b}|\ \eta\ \max_{i\in[n],\ t\in[\widehat{m}]}|\widehat{v}_{i}(t)|\leq m\eta\ \max_{i\in[n],\ t\in[\widehat{m}]}|\widehat{v}_{i}(t)|=\frac{\lambda}{2n}$ (14)
The last equality follows from the definition of $\eta$ (equation (4)). Since equation (14) holds for each $b\in[n]\setminus\mathrm{supp}(w)$, inequality (12) follows.
Establishing Inequality (13). To lower bound $p(A_{h})$, we first note that the price of item $(m+1)\in A_{h}$ satisfies $p_{m+1}=\frac{\lambda}{2}\ \max\limits_{i\in[n]}(w_{i}+\eta)\geq\frac{\lambda}{2n}$; here, the last inequality follows from the fact that $\max_{i\in[n]}w_{i}\geq 1/n$.
Now, recall that in the instance $\widehat{\mathcal{I}}=\langle[n],[\widehat{m}],\{\widehat{v}_{i}\}_{i=1}^{n}\rangle$ the set of items $\widehat{m}$] can be partitioned into three disjoint sets: goods $G$, chores $C$, and non-negatively valued items $G_{0}$ (Lemma 2). We will next show that, in allocation $\mathcal{A}$, no chore is assigned to agent $h$. Since only chores have negative prices (Proposition 2), this will give us inequality (13): $p(A_{h})\geq p_{m+1}\geq\frac{\lambda}{2n}$.
Specifically, for any chore $c\in C$ and agent $b\in[n]\setminus\mathrm{supp}(w)$, we will prove
$\eta\widehat{v}_{b}(c)>(w_{h}+\eta)\widehat{v}_{h}(c)$ (15)
Since, in allocation $\mathcal{A}$, chore $c$ must be assigned to an agent in $\operatorname*{arg\,max}_{i\in[n]}(w_{i}+\eta)\widehat{v}_{i}(c)$, inequality (15) implies that $c$ is not assigned to $h$ in $\mathcal{A}$.
Here, since $c$ is a chore, $\widehat{v}_{h}(c)<0$ and $v_{h}(c)<0$ (see Proposition 1). Moreover, $|\widehat{v}_{h}(c)|=\left|v_{h}(c)-\varepsilon_{h,c}\right|\geq\left|v_{h}(c)\right|$; the last inequality holds since $v_{h}(c)<0$ and $\varepsilon_{h,c}\geq 0$. Further, using the definition of $\lambda$ we get $|\widehat{v}_{h}(c)|\geq|v_{h}(c)|\geq\lambda$. This bound implies
$(w_{h}+\eta)|\widehat{v}_{h}(c)|$ $\geq(w_{h}+\eta)\lambda$
$\geq\frac{\lambda}{n}$ ($w_{h}\geq 1/n$ and $\eta>0$)
$>\eta\max_{i\in[n],\ t\in[\widehat{m}]}|\widehat{v}_{i}(t)|$ (equation (4))
$\geq\eta|\widehat{v}_{b}(c)|$ (16)
Since $c$ is a chore, both $\widehat{v}_{b}(c)<0$ and $\widehat{v}_{h}(c)<0$. Hence, equation (16) reduces to the stated strict inequality (15). That is, no chore $c$ is assigned to agent $h$ and, as stated, we obtain inequality (13): $p(A_{h})\geq p_{m+1}\geq\frac{\lambda}{2n}$.
In summary, we have inequalities (12) and (13). As mentioned above, these bounds imply that there exists an agent $i^{\prime}\in\mathrm{supp}(w)$ with a maximum-price bundle, $i^{\prime}\in\operatorname*{arg\,max}_{i\in[n]}p(A_{i})$. Overall, for this agent $i^{\prime}\in\mathrm{supp}(w)$, we have $w\in C_{i^{\prime}}$ and, hence, the KKM covering condition holds in the current case as well. The completes the proof of the lemma. ∎
Lemmas 8 and 9 enable us to invoke the KKM Theorem with sets $C_{i}$s as defined in Equation 10. This invocation proves that there necessarily exists a weight vector $w^{*}$ with the property that, for each agent $i\in[n]$, we have a weighted welfare-maximizing allocation $\mathcal{A}^{(i)}=(A^{(i)}_{1},\ldots,A^{(i)}_{n})$ wherein agent $i$ is price envy-free. Formally,
Theorem 10.
There exists a vector $w^{*}\in\Delta_{n-1}$ with the property that, for each $i\in[n]$, we have an allocation $\mathcal{A}^{(i)}\in\mathrm{OPT}(\textbf{{LP}}(w^{*},\eta))$ wherein agent $i\in\operatorname*{arg\,max}\limits_{j\in[n]}\ p^{*}\left(A^{(i)}_{j}\right)$; here $p^{*}=\mathrm{OPT}(\textbf{{Dual-LP}}(w^{*},\eta))$.
Proof.
Consider the sets $C_{1},\ldots,C_{n}\subseteq\Delta_{n-1}$ defined in Equation 10. Lemma 8 ensures that $C_{1},\ldots,C_{n}$ are closed. In addition, via Lemma 9, we have that these sets satisfy the KKM covering condition. Hence, invoking the KKM Theorem (Theorem 7), we obtain that there a exists a vector $w^{*}\in\cap_{i=1}^{n}C_{i}$. Write $p^{*}=\mathrm{OPT}(\textbf{{Dual-LP}}(w^{*},\eta))$. For each $i\in[n]$, we have $w^{*}\in C_{i}$. Hence, for each $i\in[n]$, by the definition of $C_{i}$, we have an allocation $\mathcal{A}^{(i)}\in\mathrm{OPT}(\textbf{{LP}}(w^{*},\eta))$ wherein agent $i\in\operatorname*{arg\,max}_{j\in[n]}{p^{*}\left(A^{(i)}_{j}\right)}$. The theorem stands proved. ∎
6 The Augmenting Tree Algorithm
This section proves the fairness and efficiency guarantee, as stated in Theorem 6, for the non-degenerate instance $\widehat{\mathcal{I}}$. Towards this, we start with weight vector $w^{*}$ and the corresponding allocations $\mathcal{A}^{(1)},\ldots,\mathcal{A}^{(n)}$, whose existence was established in Theorem 10. Then, we constructively prove that the existence of such a collection of allocations (which are price envy-free for the $n$ agents, respectively) implies the desired existence of a single allocation $\mathcal{A}^{*}$ that is $\mathrm{IEF}1$ and welfare-maximizing in $\widehat{\mathcal{I}}$. Here, our constructive proof relies on a novel Augmenting Tree Algorithm.
6.1 Price Graph
Our subsequent analysis is based on a bipartite subgraph $H$ which we define considering the weight vector $w^{*}\in\Delta_{n-1}$, identified in Theorem 10, and the associated optimal price vector $p^{*}=\mathrm{OPT}(\textbf{{Dual-LP}}(w^{*},\eta))$.
We first consider, for each agent $i\in[n]$, the items $s$ for which $i$ is the unique agent in $\operatorname*{arg\,max}_{j\in[n]}\ (w^{*}_{j}+\eta)\widehat{v}_{j}(s)$. In particular, write bundle $L_{i}\coloneqq\left\{s\in[\widehat{m}]\mid(w^{*}_{i}+\eta)\widehat{v}_{i}(s)>(w^{*}_{j}+\eta)\widehat{v}_{j}(s)\text{ for all }j\neq i\right\}$, for each $i\in[n]$. [^9] The set $L_{i}$ can be empty. Write $\mathcal{L}=(L_{1},\ldots,L_{n})$ to denote the partial allocation obtained via these bundles.
Recall that in any welfare-maximizing allocation $\mathcal{A}\in\mathrm{OPT}(\textbf{{LP}}(w^{*},\eta))$, each item $t\in[m]$ is assigned to an agent in $\operatorname*{arg\,max}_{j\in[n]}\ (w^{*}_{j}+\eta)\widehat{v}_{j}(t)$. Hence, in any welfare-maximizing allocation $\mathcal{A}=(A_{1},\ldots,A_{n})$ the containment $L_{i}\subseteq A_{i}$ holds for each $i\in[n]$.
We define the bipartite graph $H=([n]\cup U,E)$ with the left part as the set of agents, $[n]$, and the right part $U\coloneqq[\widehat{m}]\setminus\left(\cup_{i=1}^{n}L_{i}\right)$. The definition of $L_{i}$s implies that the right part $U$ consists of the items $t$ with $\left|\operatorname*{arg\,max}_{j\in[n]}\ (w^{*}_{j}+\eta)\widehat{v}_{j}(t)\right|\geq 2$. Moreover, we include $(i,t)\in[n]\times U$ in the edge set $E$ iff $i\in\operatorname*{arg\,max}_{j\in[n]}(w^{*}_{j}+\eta)\widehat{v}_{j}(t)$.
We will use $\Gamma(x)$ to denote the vertices adjacent to each $x\in[n]\cup U$ in $H=([n]\cup U,E)$. Note that the construction of the edge set $E$ and the definition of $U$ ensure that each item $t\in U$ has degree at least two in the bipartite graph, $|\Gamma(t)|\geq 2$.
The lemma below provides a key property of $H$: it is acyclic.
Lemma 11.
The bipartite graph $H=([n]\cup U,E)$ is a forest.
Proof.
First, we note that in the bipartite graph $H=([n]\cup U,E)$, for every edge $(i,t)\in E$ we have $\widehat{v}_{i}(t)\neq 0$: The optimal prices $p^{*}_{t}\neq 0$ for each item $t\in[\widehat{m}$] (Proposition 2). In addition, by construction of the edge set $E$, for every $(i,t)\in E$, we have $p^{*}_{t}=(w^{*}_{i}+\eta)\widehat{v}_{i}(t)\neq 0$. Since $(w^{*}_{i}+\eta)>0$, it holds that $\widehat{v}_{i}(t)\neq 0$.
Assume, towards a contradiction, that $H$ admits a simple cycle $C=(i_{1},t_{1},i_{2},t_{2},\ldots,i_{k},t_{k},i_{1})$, with $i_{1},\ldots,i_{k}\in[n]$ and $t_{1},\ldots,t_{k}\in U$. Here, each $(i_{\ell},t_{\ell}),(t_{\ell},i_{\ell+1})\in E$ and, hence, as noted above, the values $\widehat{v}_{i_{\ell}}(t_{\ell}),\widehat{v}_{i_{\ell+1}}(t_{\ell})\neq 0$, for each $\ell\in[k]$. [^10] By convention we set $i_{k+1}=i_{1}$. Therefore, the product $\prod_{\ell=1}^{k}\left(\frac{\widehat{v}_{i_{\ell+1}}(t_{\ell})}{\widehat{v}_{i_{\ell}}(t_{\ell})}\right)$ is nonzero and satisfies
$\prod_{\ell=1}^{k}\left(\frac{\widehat{v}_{i_{\ell+1}}(t_{\ell})}{\widehat{v}_{i_{\ell}}(t_{\ell})}\right)=\prod_{\ell=1}^{k}\left(\frac{(w^{*}_{i_{\ell+1}}+\eta)\widehat{v}_{i_{\ell+1}}(t_{\ell})}{(w^{*}_{i_{\ell}}+\eta)\widehat{v}_{i_{\ell}}(t_{\ell})}\right)$ (17)
The last equality is obtained by multiplying and dividing by the positive terms $\{(w^{*}_{i_{\ell}}+\eta)\}_{\ell\in[k]}$ to the left-hand-side of the equation, and a telescoping shift.
Since both $(i_{\ell},t_{\ell}),(t_{\ell},i_{\ell+1})\in E$, we have $p^{*}_{t_{\ell}}=(w^{*}_{i_{\ell+1}}+\eta)\widehat{v}_{i_{\ell+1}}(t_{\ell})=(w^{*}_{i_{\ell}}+\eta)\widehat{v}_{i_{\ell}}(t_{\ell})$. That is, each ratio $\left(\frac{(w^{*}_{i_{\ell+1}}+\eta)\widehat{v}_{i_{\ell+1}}(t_{\ell})}{(w^{*}_{i_{\ell}}+\eta)\widehat{v}_{i_{\ell}}(t_{\ell})}\right)=1$ and, hence, equation (17) reduces to $\prod_{\ell=1}^{k}\left(\frac{\widehat{v}_{i_{\ell+1}}(t_{\ell})}{\widehat{v}_{i_{\ell}}(t_{\ell})}\right)=1$.
This, however, contradicts Lemma 3, since $C$ is a cycle in the complete bipartite graph $K_{n,\widehat{m}}$ as well. Hence, by way of contradiction, we obtain that $H$ is acyclic, i.e., a forest. The lemma stands proved. ∎
The following lemma provides another useful property of the bipartite graph $H$: Any completion of the partial allocation $\mathcal{L}=(L_{1},\ldots,L_{n})$ by assigning the items in $U=[\widehat{m}]\setminus\left(\cup_{i\in[n]}\ L_{i}\right)$ along the edges in $H$ provides a welfare-maximizing allocation.
Specifically, define the collection of (complete) allocations
$\mathcal{O}^{*}\coloneqq\left\{(B_{1},\ldots,B_{n})\in\Pi_{n}([\widehat{m}])\mid L_{i}\subseteq B_{i}\subseteq L_{i}\cup\Gamma(i)\text{ for each }i\in[n]\right\}$ (18)
We next show that $\mathcal{O}^{*}$ is same as $\mathrm{OPT}(\textbf{{LP}}(w^{*},\eta))$, the set of welfare-maximizing allocations.
Lemma 12.
Any allocation $\mathcal{B}\in\mathcal{O}^{*}$ if and only if $\mathcal{B}\in\mathrm{OPT}(\textbf{{LP}}(w^{*},\eta))$.
Proof.
If $\mathcal{B}=(B_{1},\ldots,B_{n})\in\mathcal{O}^{*}$, then, by definition, for all $i\in[n]$, we have $L_{i}\subseteq B_{i}\subseteq L_{i}\cup\Gamma(i)$. Hence, if item $t\in B_{i}$, then $(i,t)\in E$, i.e., $p^{*}_{t}=(w^{*}_{i}+\eta)\widehat{v}_{i}(t)$. This implies that the objective value of $\textbf{{LP}}(w^{*},\eta)$ for the allocation $\mathcal{B}$ is $\sum_{i\in[n]}\sum_{t\in B_{i}}(w^{*}_{i}+\eta)\widehat{v}_{i}(t)=\sum_{t\in[\widehat{m}]}p^{*}_{t}$. Note that $\sum_{t\in[\widehat{m}]}p^{*}_{t}$ is also the objective value of $\textbf{{Dual-LP}}(w^{*},\eta)$. Hence, by strong duality, we get that $\mathcal{B}\in\mathrm{OPT}(\textbf{{LP}}(w^{*},\eta))$.
To prove the other direction, consider any $\mathcal{B}^{\prime}\in\mathrm{OPT}(\textbf{{LP}}(w^{*},\eta))$. Optimality of $\mathcal{B}^{\prime}=(B^{\prime}_{1},\ldots,B^{\prime}_{n})$ implies that if $t\in B^{\prime}_{i}$, then $(w^{*}_{i}+\eta)\widehat{v}_{i}(t)=\max_{j\in[n]}\ (w^{*}_{j}+\eta)\widehat{v}_{j}(t)$.
For item $t$, if $i$ is the unique agent in $\operatorname*{arg\,max}_{j\in[n]}\ (w^{*}_{j}+\eta)\widehat{v}_{j}(t)$, then $t\in L_{i}$. Otherwise, by construction of the edge set $E$ in the bipartite graph $H$, we have $(i,t)\in E$. Hence, for each $i\in[n]$, we have $L_{i}\subseteq A_{i}\subseteq L_{i}\cup\Gamma(i)$, implying that $\mathcal{B}^{\prime}\in\mathcal{O}^{*}$. This completes the proof of the lemma. ∎
6.2 Min-Max Allocations
A key construct in our analysis is the min-max price threshold $\tau\in\mathbb{R}$, defined as
$\tau\coloneqq\min_{(B_{1},\ldots,B_{n})\in\mathcal{O}^{*}}\ \ \max_{j\in[n]}\ p^{*}(B_{j})$ (19)
In addition, write $\mathcal{M}^{*}$ to denote all the min-max allocations
$\mathcal{M}^{*}=\left\{(B_{1},\ldots,B_{n})\in\mathcal{O}^{*}\mid\max_{j\in[n]}\ p^{*}(B_{j})=\tau\right\}$ (20)
By definition, $\mathcal{M}^{*}\subseteq\mathcal{O}^{*}$.
Additionally, for each agent $i\in[n]$ and any subset of items $S$, we define $p^{+}_{i}(S)$ as the maximum price that can be achieved by toggling (adding or removing) at most one item from $\Gamma(i)$ within SS. [^11] Recall that $\Gamma(i)$ denotes the set of neighbors of $i$ in the bipartite graph $H$. Also, if item $t\in S$, then it holds that $S\triangle\{t\}=S\setminus\{t\}$. Otherwise, for $t\notin S$, we have $S\triangle\{t\}=S\cup\{t\}$.
$p^{+}_{i}(S)\coloneqq\max\limits_{\begin{subarray}{c}T\subseteq\Gamma(i),\\ |T|\leq 1\end{subarray}}\ p^{*}(S\triangle T)$ (21)
[…]
We will work with a potential $\varphi^{+}(\cdot)$ that, for any allocation $\mathcal{B}=(B_{1},\ldots,B_{n})$, counts the number of agents whose bundle prices can be toggled to at least $\tau$.
$\varphi^{+}(\mathcal{B})=\Big|\left\{i\in[n]\mid p^{+}_{i}(B_{i})\geq\tau\right\}\Big|$ (22)
For any min-max allocation $\mathcal{B}\in\mathcal{M}^{*}$ (since $\max_{j\in[n]}p^{*}(B_{j})=\tau$), we have $\varphi^{+}(\mathcal{B})\geq 1$.
The following theorem is a key result of this section, and it shows that there exists a min-max allocation $\mathcal{A}^{*}$ wherein the bundle prices of all the agents can be toggled to $\tau$, or above. We provide a constructive proof of Theorem 13 in Section 6.4.
[…]
Theorem 13.
There exists an allocation $\mathcal{A}^{*}\in\mathcal{M}^{*}$ such that $\varphi^{+}(\mathcal{A}^{*})=n$.
6.3 Proof of Theorem 6
See 6
Proof.
Let $w^{*}\in\Delta_{n-1}$ be the weight vector whose existence is guaranteed by Theorem 10, and $p^{*}=\mathrm{OPT}(\textbf{{Dual-LP}}(w^{*},\eta))$ be the associated optimal prices. Further, the collection of allocations $\mathcal{O}^{*}$ and $\mathcal{M}^{*}$ are as defined in equations (18) and (20), respectively. We will consider the allocation $\mathcal{A}^{*}\in\mathcal{M}^{*}$ provided in Theorem 13.
Since allocation $\mathcal{A}^{*}\in\mathcal{M}^{*}\subseteq\mathcal{O}^{*}$, Lemma 12 implies that, as desired, allocation $\mathcal{A}^{*}$ is welfare maximizing, $\mathcal{A}^{*}\in\mathrm{OPT}(\textbf{{LP}}(w^{*},\eta))$.
We will complete the proof by proving next that $\mathcal{A}^{*}=(A^{*}_{1},\ldots,A^{*}_{n})$ is $\mathrm{IEF}1$ in the instance $\widehat{\mathcal{I}}$. For allocation $\mathcal{A}^{*}$ we have $\varphi^{+}(\mathcal{A}^{*})=n$. Hence, for each agent $i\in[n]$, it holds that
$p^{+}_{i}(A^{*}_{i})\geq\tau=\max_{j\in[n]}p^{*}(A^{*}_{j})$ (23)
[…]
$p^{*}(A^{*}_{i}\triangle T)\geq p^{*}(A^{*}_{j})$ (24)
Since $\mathcal{A}^{*}\in\mathcal{O}^{*}$, we have $L_{i}\subseteq A^{*}_{i}\subseteq L_{i}\cup\Gamma(i)$; see equation (18). Also, $T\subseteq\Gamma(i)$. Hence, we have $A^{*}_{i}\triangle T\subseteq L_{i}\cup\Gamma(i)$. Note that for each item $t\in L_{i}\cup\Gamma(i)$ and, hence, for each $t\in A^{*}_{i}\triangle T$ it holds that $(w^{*}_{i}+\eta)\ \widehat{v}_{i}(t)=p^{*}_{t}$. Using this equality for all the items in $A^{*}_{i}\triangle T$, we extend equation (24) to
$(w^{*}_{i}+\eta)\ \widehat{v}_{i}(A^{*}_{i}\triangle T)=p^{*}(A^{*}_{i}\triangle T)\geq p^{*}(A^{*}_{j})\geq(w^{*}_{i}+\eta)\ \widehat{v}_{i}(A_{j}^{*})$ (25)
The last inequality follows from the fact that, for each item $s\in[\widehat{m}$], the optimal price $p^{*}_{s}=\max_{k\in[n]}\ (w^{*}_{k}+\eta)\widehat{v}_{k}(s)\geq(w^{*}_{i}+\eta)\widehat{v}_{i}(s)$.
Since $(w^{*}_{i}+\eta)>0$, equation (25) gives us $\widehat{v}_{i}(A^{*}_{i}\triangle T)\geq\widehat{v}_{i}(A^{*}_{j})$, for each $j\in[n]$. This bound implies that $\mathcal{A}^{*}$ is $\mathrm{IEF}1$ in $\widehat{\mathcal{I}}$, since for each agent $i\in[n]$, there exists a subset $T\subseteq[\widehat{m}$]—of size at most one—such that $\widehat{v}_{i}(A^{*}_{i}\triangle T)\geq\max_{j\in[n]}\widehat{v}_{i}(A^{*}_{j})$. The theorem stands proved. ∎
6.4 Proof of Theorem 13
[…]
We begin by describing the AugmentTree Algorithm (Algorithm 1). AugmentTree Algorithm starts with input allocation $\mathcal{S}=(S_{1},\ldots,S_{n})\in\mathcal{M}^{*}$ for which $\varphi^{+}(\mathcal{S})<n$. Hence, there exists an agent $r\in[n]$ such that $p^{+}_{r}(S_{r})<\tau$. In the bipartite $H=([n]\cup U,E)$, we consider the connected component that contains $r$. Given that $H$ is a forest (Lemma 11), the identified connect component is a tree, and we root this tree at $r$ itself. We will, throughout, write $R$ to denote this rooted tree.
Also, for each node $x$ in the rooted tree $R$, write $\mathrm{par}(x)\in\Gamma(x)$ to denote the (unique) parent of $x$. Since $H$ is bipartite between agents and items, for each agent $i\in[n]$ in the tree, the parent $\mathrm{par}(i)$ is an item. Conversely, for any item $t$ in the tree, $\mathrm{par}(t)$ is an agent. Also, for the root $r$ we set $\mathrm{par}(r)=\emptyset$.
Furthermore, let $\mathrm{ch}(x)=\Gamma(x)\setminus\{\mathrm{par}(x)\}$ denote the children of node $x$ in the rooted tree $R$. Also, as mentioned previously, each item $t$ in the bipartite graph $H$ (and, hence, in $R$) has degree at least two. Therefore, for each item $t$ in the tree, we have $\mathrm{ch}(t)\neq\emptyset$.
Finally, for each agent $i$ in the rooted tree, we will use $\overline{\mathrm{ch}}(i)$ to denote the agents that are immediate descendants of $i$, i.e., $\overline{\mathrm{ch}}(i)\coloneqq\bigcup_{t\in\mathrm{ch}(i)}\ \mathrm{ch}(t)$. [^12] If, for any agent $i$, the set $\mathrm{ch}(i)$ is empty, then so is $\overline{\mathrm{ch}}(i)$.
The following proposition will be used in the analysis of the algorithm – it directly follows from the fact that, in a rooted tree, no two distinct nodes can have a common child. [^13] Indeed, two distinct nodes can have a common parent.
Proposition 3.
Let $i\in[n]$ be an agent in the rooted tree $R$, and let item $t\in\mathrm{ch}(i)$. If $t$ is a neighbor of an agent $a\in[n]$ (i.e., $t\in\Gamma(a)$), then necessarily $\mathrm{par}(a)=t$ and $a\in\overline{\mathrm{ch}}(i)$.
[…]
If, after this exchange, an agent $a\in\overline{\mathrm{ch}}(i)$ is left with a bundle satisfying $p^{+}_{a}(\widehat{S}_{a})<\tau$, then we add $a$ to $\mathcal{Q}$ (Lines 1), in order to process it subsequently and ensure that the final bundle of $a$ satisfies $p^{+}_{a}(\widehat{S}_{a})\geq\tau$. Proceeding in this manner, the algorithm iterates as long as $\mathcal{Q}\neq\emptyset$. We prove that the algorithm terminates and always returns an allocation $\widehat{\mathcal{S}}$ such that $\varphi^{+}(\widehat{\mathcal{S}})>\varphi^{+}(\mathcal{S})$. In addition, the algorithm maintains throughout that $\max_{j\in[n]}p^{*}(\widehat{S}_{j})\leq\tau$, i.e., even the returned allocation $\widehat{\mathcal{S}}\in\mathcal{M}^{*}$. As mentioned previously, this guaranteed success of the algorithm proves Theorem 13.
1
2
3
Input : Allocation $\mathcal{S}=(S_{1},\dots,S_{n})\in\mathcal{M}^{*}$ (i.e., $\max_{j}p^{*}(S_{j})=\tau$) with $\varphi^{+}(\mathcal{S})<n$.
Output : Allocation $\widehat{\mathcal{S}}=(\widehat{S}_{1},\dots,\widehat{S}_{n})\in\mathcal{M}^{*}$ satisfying $\varphi^{+}(\widehat{\mathcal{S}})>\varphi^{+}(\mathcal{S})$.
4
5Let $r\in[n]$ to be any agent with $p^{+}_{r}(S_{r})<\tau$. Initialize set $\mathcal{Q}=\{r\}$.
6
7As defined above, $R$ is the tree rooted at $r$ and, for each node $x$ in the tree, $\mathrm{par}(x)$ denotes the parent of $x$. In addition, $\mathrm{ch}(x)=\Gamma(x)\setminus\{\mathrm{par}(x)\}$ are the children of $x$. Also, for each agent $i$ in the tree, $\overline{\mathrm{ch}}(i)=\cup_{t\in\mathrm{ch}(i)}\ \mathrm{ch}(t)$ denotes the agents that are immediate descendants of $i$.
8 Initialize $\widehat{S}_{j}=S_{j}$ for all agents $j\in[n]$.
9
101exwhile $\mathcal{Q}\neq\emptyset$ do
11   Let $i\in\mathcal{Q}$ be any agent in $\mathcal{Q}$. Update $\mathcal{Q}\leftarrow\mathcal{Q}\setminus\{i\}$.
12
13  1exSet $X\subseteq\mathrm{ch}(i)$ to be an item subset that satisfies: (i) $p^{+}_{i}(\widehat{S}_{i}\triangle X)\geq\tau$, (ii) $p^{*}(\widehat{S}_{i}\triangle X)\leq\tau$, and (iii) $p^{*}(\widehat{S}_{i}\triangle\{t\})>p^{*}(\widehat{S}_{i})$ for each $t\in X$. {Existence of $X$ is guaranteed via Lemma 14.}
14
15  1exfor all items $t\in X$ do
16      If item $t\in\widehat{S}_{i}$, then set $a$ to be any agent in $\mathrm{ch}(t)\neq\emptyset$. Otherwise, if $t\notin\widehat{S}_{i}$, then let $a\in\mathrm{ch}(t)$ be the agent such that $t\in\widehat{S}_{a}$. {We prove that such an agent $a\in\overline{\mathrm{ch}}(i)$ exists.}
17      Update $\widehat{S}_{a}\leftarrow\widehat{S}_{a}\triangle\{t\}$.
18
19     1exIf $p^{+}_{a}(\widehat{S}_{a})<\tau$, then include $a$ for processing: $\mathcal{Q}\leftarrow\mathcal{Q}\cup\{a\}$. {Otherwise, if $p^{+}_{a}(\widehat{S}_{a})\geq\tau$, then we do not include $a$ in $\mathcal{Q}$.}
20
21
22  1exUpdate $\widehat{S}_{i}\leftarrow\widehat{S}_{i}\triangle X$
23
24return Allocation $\widehat{\mathcal{S}}=(\widehat{S}_{1},\ldots,\widehat{S}_{n})$.
Algorithm 1 AugmentTree Algorithm
Before analyzing the algorithm, we prove a key lemma that encapsulates the existential guarantee of Theorem 10.
Lemma 14.
For any agent $i\in[n]$ in the rooted tree, let $S$ be an item subset such that $L_{i}\subseteq S\subseteq L_{i}\cup\Gamma(i)$ and $p^{+}_{i}(S)<\tau$. Then, there exists subset $X\subseteq\mathrm{ch}(i)$ satisfying the following properties
  1. (i)
    $p^{+}_{i}(S\triangle X)\geq\tau$,
  2. (ii)
    $p^{*}(S\triangle X)\leq\tau$, and
  3. (iii)
    $p^{*}(S\triangle\{t\})>p^{*}(S)$ for each $t\in X$.
Proof.
For the agent $i\in[n]$, consider the allocation $\mathcal{A}^{(i)}\in\mathrm{OPT}(\textbf{{LP}}(w^{*},\eta))$ from Theorem 10, in which agent $i$ is price envy-free; in particular, $p^{*}(A^{(i)}_{i})=\max_{j\in[n]}\ p^{*}(A^{(i)}_{j})$.
Lemma 12 implies $\mathcal{A}^{(i)}\in\mathcal{O}^{*}$ and, hence, we have $L_{i}\subseteq A^{(i)}_{i}\subseteq L_{i}\cup\Gamma(i)$; see equation (18). Furthermore, the definition of $\tau$ (equation (19)) gives us $p^{*}(A^{(i)}_{i})=\max_{j\in[n]}\ p^{*}(A^{(i)}_{j})\geq\tau$.
Now, consider the set
$K\coloneqq\left\{t\in A^{(i)}_{i}\triangle S\mid p^{*}(S\triangle\{t\})>p^{*}(S)\right\}$ (26)
The set $K$ consists of the chores in $S\setminus A^{(i)}_{i}$ and the goods (along with non-negatively valued items) in $A^{(i)}_{i}\setminus S$. Moreover, note that the bound $p^{*}\left(A^{(i)}_{i}\right)\geq\tau$ implies that for the set $K$ we have
$p^{*}(S\triangle K)\geq\tau$ (27)
Further, by the lemma assumption, the given set $S$ satisfies $p^{+}_{i}(S)<\tau$ and, hence, $p^{*}(S)<\tau$. This bound and inequality (27) ensure that $K\neq\emptyset$. We define $\widetilde{K}\subseteq K$ to be a minimal subset of $K$ with the property that $p^{*}\left(S\triangle\widetilde{K}\right)\geq\tau$. As in the case of $K$, it holds that $\widetilde{K}\neq\emptyset$. Finally, we obtain the desired subset $X$ as follows
[…]
We first show that $X\subseteq\mathrm{ch}(i)$. Towards this, note the containments $L_{i}\subseteq A^{(i)}_{i}\subseteq L_{i}\cup\Gamma(i)$ along with $L_{i}\subseteq S\subseteq L_{i}\cup\Gamma(i)$. Hence, $\Gamma(i)\supseteq A^{(i)}_{i}\triangle S\supseteq K\supseteq\widetilde{K}$. By construction of $X$, we have $\mathrm{par}(i)\notin X$. Therefore, $X\subseteq\Gamma(i)\setminus\{\mathrm{par}(i)\}=\mathrm{ch}(i)$.
We will complete the proof by establishing that $X$ satisfies the three properties stated in the lemma. First, by the minimality of $\widetilde{K}$, we have $p^{*}(S\triangle X)<\tau$; this establishes property (ii) from the lemma.
Next, note that $X$ is obtained by removing one item from $\widetilde{K}\subseteq\Gamma(i)$ and, hence, $p^{+}_{i}(S\triangle X)\geq p^{*}\left(S\triangle\widetilde{K}\right)\geq\tau$. That is, property (i) holds as well.
For property (iii), recall that, by definition, each item $t\in K$ satisfies $p^{*}(S\triangle\{t\})>p^{*}(S)$. Since $X\subseteq\widetilde{K}\subseteq K$, the property follows. This completes the proof of the lemma. ∎
We analyze AugmentTree Algorithm by showing that each iteration of its while-loop executes successfully (Lemma 15) and that the while-loop necessarily terminates (Corollary 17). Finally, we establish Theorem 13 by proving that the algorithm always returns a complete allocation $\widehat{\mathcal{S}}\in\mathcal{M}^{*}$ with an increased $\varphi^{+}(\cdot)$ value.
Lemma 15.
Each iteration of the while-loop in the AugmentTree Algorithm executes successfully and maintains an allocation $\widehat{\mathcal{S}}\in\mathcal{M}^{*}$. Furthermore, at the end of the iteration (Line 1), for the agent $i$ selected in that iteration, it holds that $p^{+}_{i}(\widehat{S}_{i})\geq\tau$.
Proof.
Fix any iteration of the while-loop, and let $i\in\mathcal{Q}$ be the agent under consideration. At the beginning of this iteration, for the current allocation $\widehat{\mathcal{S}}=(\widehat{S}_{1},\ldots,\widehat{S}_{n})$, we have $p^{+}_{i}(\widehat{S}_{i})<\tau$. This follows from the fact that any agent $i$ (including the root $r$) is added to $\mathcal{Q}$ only when $p^{+}_{i}(\widehat{S}_{i})<\tau$. In addition, inductively, we have that the current allocation $\widehat{\mathcal{S}}\in\mathcal{M}^{*}\subseteq\mathcal{O}^{*}$. Hence, $L_{i}\subseteq\widehat{S}_{i}\subseteq L_{i}\cup\Gamma(i)$. These observations allow us to apply Lemma 14 with $S=\widehat{S}_{i}$ and obtain an item subset $X\subseteq\mathrm{ch}(i)$ that satisfies the requirements specified in Line 1.
Next, we observe that, for each item $t\in X\subseteq\mathrm{ch}(i)$, the desired agent $a$ (Line 1) exists: Recall that $\mathrm{ch}(s)\neq\emptyset$ for any item $s$. Also, the current allocation $\widehat{\mathcal{S}}\in\mathcal{M}^{*}\subseteq\mathcal{O}^{*}$ and, hence, item $t$ must to assigned (in $\widehat{\mathcal{S}}$) to an agent in $\Gamma(t)$. Therefore, if $t\in\widehat{S}_{i}$, then (as stated in Line 1) we can select any agent $a\in\mathrm{ch}(t)$. Otherwise, if $t\notin\widehat{S}_{i}$, it must be the case that $t\in\widehat{S}_{a}$ for an agent $a\in\mathrm{ch}(t)$ that satisfies $\mathrm{par}(a)=t$ and $a\in\overline{\mathrm{ch}}(i)$; see Proposition 3.
The existence of subset $X$ and agents $a$ ensure that the iteration executes successfully. Further, for the selected subset $X$ (Lemma 14) we have $p^{+}_{i}(\widehat{S}_{i}\triangle X)\geq\tau$. Hence, at the end of the iteration, i.e., after the update $\widehat{S}_{i}\leftarrow\widehat{S}_{i}\triangle X$ (in Line 1), we have $p^{+}_{i}(\widehat{S}_{i})\geq\tau$, as stated.
It remains to show that, even after the updates in Lines 1 and 1, the allocation $\widehat{\mathcal{S}}\in\mathcal{M}^{*}$. Note that the algorithm maintains the invariant that each item $t$ is assigned to an agent in $\Gamma(t)$. That is, the containments $L_{j}\subseteq\widehat{S}_{j}\subseteq L_{j}\cup\Gamma(j)$ continue to hold and, hence, the updated allocation satisfies $\widehat{\mathcal{S}}\in\mathcal{O}^{*}$. Here, for agent $i$, property (ii) of subset $X$ ensures that for the updated bundle $\widehat{S}_{i}$ we have $p^{*}(\widehat{S}_{i})\leq\tau$. In addition, property (iii) ensures that for agents $a$ considered in the iteration $p^{*}(\widehat{S}_{a})$ in fact decreases after the update in Line 1: $p^{*}(\widehat{S}_{i})+p^{*}(\widehat{S}_{a})=p^{*}(\widehat{S}_{i}\triangle\{t\})+p^{*}(\widehat{S}_{a}\triangle\{t\})$ and $p^{*}(\widehat{S}_{i}\triangle\{t\})>p^{*}(\widehat{S}_{i})$. Therefore, for the allocation $\widehat{\mathcal{S}}$ maintained at the end of iteration, we have $p^{*}(\widehat{S}_{j})\leq\tau$, for each agent $j$, i.e., as desired $\widehat{\mathcal{S}}\in\mathcal{M}^{*}$. This completes the proof of the lemma. ∎
Lemma 16.
Any agent $i\in[n]$ is included at most once in the set $\mathcal{Q}$ throughout the execution of AugmentTree Algorithm.
Proof.
The root agent $r$ is included in $\mathcal{Q}$ at initialization. Any other agent $j\in[n]$ is included in $\mathcal{Q}$ only when its immediate ancestor is being processed in the while-loop iteration. In particular, if $j$ is added to $\mathcal{Q}$ in some iteration, then the agent $i$ being processed at that time must satisfy $j\in\overline{\mathrm{ch}}(i)$ (see Lines 1 and 1). Since the root $r$ has no ancestors, it is never included in $\mathcal{Q}$—and, hence, never processed—a second time.
By inductively applying this observation down the levels of the rooted tree, we obtain that no agent is added to $\mathcal{Q}$ more than once. The lemma stands proved. ∎
Corollary 17.
The AugmentTree Algorithm (Algorithm 1) always terminates and returns a complete allocation $\widehat{\mathcal{S}}=(\widehat{S}_{1},\ldots,\widehat{S}_{n})\in\mathcal{M}^{*}$.
Proof.
Since each iteration of the while-loop executes successfully (Lemma 15) and the while-loop can iterate at most $n$ times (Lemma 16), it follows that the algorithm necessarily terminates. Furthermore, given that the input allocation is contained in $\mathcal{M}^{*}$ and the algorithm maintains this containment, $\widehat{\mathcal{S}}\in\mathcal{M}^{*}$ (Lemma 15), we obtain that the returned allocation belongs to $\mathcal{M}^{*}$ as well. This completes the proof of the corollary. ∎
We now establish Theorem 13.
Proof of Theorem 13.
As mentioned previously, to prove the theorem it suffices to show that, for any input allocation $\mathcal{S}\in\mathcal{M}^{*}$, AugmentTree Algorithm returns another allocation $\widehat{\mathcal{S}}\in\mathcal{M}^{*}$ with a strictly higher $\varphi^{+}(\cdot)$ value. In particular, by repeatedly applying the algorithm, we obtain the desired allocation $\mathcal{A}^{*}\in\mathcal{M}^{*}$ with $\varphi^{+}(\mathcal{A}^{*})=n$, thereby establishing the theorem.
Corollary 17 ensures that the algorithm successfully finds $\widehat{\mathcal{S}}\in\mathcal{M}^{*}$. We now complete the proof by showing that $\varphi^{+}(\widehat{\mathcal{S}})>\varphi^{+}(\mathcal{S})$.
[…]
This observation and the inclusion of $r$ (as an agent which satisfies $p^{+}_{r}(\widehat{S}_{r})\geq\tau$) imply that $\varphi^{+}(\widehat{\mathcal{S}})>\varphi^{+}(\mathcal{S})$. The theorem stands proved. ∎
In summary, through the chain of implications illustrated in Figure 1, we establish the existence of fair and efficient allocations for indivisible mixed manna under additive valuations.
Figure 1: Chain of implications established in Sections 4, 5, and 6.
7 $\mathrm{IEF}1$ and $\mathrm{fPO}$ for Mixed Manna
This section strengthens Theorem 1 on the efficiency front. In particular, we show that, for indivisible mixed manna with additive valuations, there always exists an (integral) allocation that is both $\mathrm{IEF}1$ and fractionally Pareto efficient ($\mathrm{fPO}$). The proof for this result relies on LP duality and limit arguments.
Theorem 18.
Every fair division instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$, with indivisible mixed manna and additive valuations, admits an $\mathrm{IEF}1$ and $\mathrm{fPO}$ allocation.
Proof.
For the given instance $\mathcal{I}$, consider the non-degenerate instance $\widehat{\mathcal{I}}$ as constructed in Section 4.1. Recall that the construction of $\widehat{\mathcal{I}}$ is applicable as long as the chosen positive parameter $\varepsilon>0$ satisfies the strict inequality in equation (1). We will denote $\widehat{\mathcal{I}}$ as $\widehat{\mathcal{I}}_{1}$.
[…]
The claim below identifies a useful subsequence of the instances $\left\{\widehat{\mathcal{I}}_{k}\right\}_{k\in\mathbb{N}}$ and the corresponding weight vectors $\left\{w_{k}\in\Delta_{n-1}\right\}_{k\in\mathbb{N}}$.
Claim 3.
[…]
  1. (1)
    For all $k\in\mathbb{N}$, allocation $\mathcal{A}^{*}$ is $\mathrm{IEF}1$ in $\widehat{\mathcal{I}}_{\ell_{k}}$ and $\mathcal{A}^{*}\in\mathrm{OPT}(\textbf{{LP}}(w_{\ell_{k}},\eta,\widehat{\mathcal{I}}_{\ell_{k}}))$, where $w_{\ell_{k}}$ is the weight vector for $\widehat{\mathcal{I}}_{\ell_{k}}$.
  2. (2)
    The sequence $\left\{w_{\ell_{k}}\right\}_{k\in\mathbb{N}}$ converges to some $w^{*}\in\Delta_{n-1}$, and
  3. (3)
    The sequence of optimal price vectors $\left\{p^{\ell_{k}}=\mathrm{OPT}({\textbf{{Dual-LP}}}(w_{{\ell_{k}}},\eta,\widehat{\mathcal{I}}_{{{\ell}_{k}}}))\right\}_{k\in\mathbb{N}}$ also converges to a point $p^{*}\in\mathbb{R}^{\widehat{m}}$.
Proof.
Consider the infinite sequence of instances $\left\{\widehat{\mathcal{I}}_{k}\right\}_{k\in\mathbb{N}}$ and the corresponding allocations $\left\{\mathcal{A}_{k}\right\}_{k\in\mathbb{N}}$ that satisfy $\mathcal{A}_{k}\in\mathrm{OPT}(\textbf{{LP}}(w_{k},\eta,\widehat{\mathcal{I}}_{k}))$. Here, $\mathcal{A}_{k}$ is $\mathrm{IEF}1$ in $\widehat{\mathcal{I}}_{k}$, for each $k$. Since the number of allocations is finite, there exists an allocation $\mathcal{A}^{*}\in\Pi_{n}([m])$ and an infinite subsequence $\left\{\widehat{\mathcal{I}}_{a_{k}}\right\}_{k\in\mathbb{N}}$ such that $\mathcal{A}_{a_{k}}=\mathcal{A}^{*}$ for each $k$.
Now, for the subsequence $\left\{\widehat{\mathcal{I}}_{a_{k}}\right\}_{k\in\mathbb{N}}$, consider the associated weights $\left\{w_{a_{k}}\right\}_{k\in\mathbb{N}}$. Since these weights lie in the bounded set $\Delta_{n-1}$, we obtain, via the Bolzano-Weierstrass theorem, a further subsequence $\{w_{\ell_{k}}\}_{k\in\mathbb{N}}$ that converges to a point $w^{*}\in\Delta_{n-1}$; here, the weight vector $w_{\ell_{k}}$ is associated with instance $\widehat{\mathcal{I}}_{\ell_{k}}$. This establishes property (2) in the lemma statement.
Note that, for each $k\in\mathbb{N}$, we have $\mathcal{A}^{*}\in\mathrm{OPT}(\textbf{{LP}}(w_{\ell_{k}},\eta,\widehat{\mathcal{I}}_{\ell_{k}}))$ and $\mathcal{A}^{*}$ is $\mathrm{IEF}1$ in $\widehat{\mathcal{I}}_{\ell_{k}}$. This gives us property (1).
Finally, for property (3), we use the fact that the optimal prices $p^{\ell_{k}}=\mathrm{OPT}({\textbf{{Dual-LP}}}(w_{{\ell_{k}}},\eta,\widehat{\mathcal{I}}_{{{\ell}_{k}}}))$ satisfy $p^{\ell_{k}}_{t}=\max\limits_{i\in[n]}\ \big((w_{\ell_{k}})_{i}+\eta\big)\widehat{v}^{\ell_{k}}_{i}(t)$ for each item $t\in[\widehat{m}$]; here $\widehat{v}^{\ell_{k}}_{i}(t)$ is the value of $t$ for agent $i$ in $\widehat{\mathcal{I}}_{{{\ell}_{k}}}$. As $k$ tends to infinity, $w_{\ell_{k}}$ tends to $w^{*}$ and $\widehat{v}^{\ell_{k}}_{i}(t)$ tends to $v_{i}(t)$, for each $i$ and $t\in[m]$. Also, $\widehat{v}^{\ell_{k}}_{i}(m+1)=\lambda/2$ for each $i\in[n]$. Since the prices are continuous functions of the converging variables, by the sequential criterion of continuity, we obtain that $\lim_{k\to\infty}p^{\ell_{k}}_{t}=\max_{i\in[n]}\ (w^{*}_{i}+\eta)v_{i}(t)$, for each $t\in[m]$, and $\lim_{k\to\infty}p^{\ell_{k}}_{\widehat{m}}=\max_{i\in[n]}\ (w^{*}_{i}+\eta)\lambda/2$, implying that $\left\{p^{\ell_{k}}\right\}_{k\in\mathbb{N}}$ converges. Write the limit point as $p^{*}\in\mathbb{R}^{\widehat{m}}$. This establishes property (3) and completes the proof of the claim. ∎
Using 3 and for the given instance $\mathcal{I}$, we will next prove that $(A^{*}_{1}\setminus\{m+1\},\ldots,A^{*}_{n}\setminus\{m+1\})\in\mathrm{OPT}(\textbf{{LP}}(w^{*},\eta,\mathcal{I}))$ and $(A^{*}_{1}\setminus\{m+1\},\ldots,A^{*}_{n}\setminus\{m+1\})$ is $\mathrm{IEF}1$ in $\mathcal{I}$. This will imply that $\mathcal{A}^{*}$ is $\mathrm{fPO}$ and $\mathrm{IEF}1$ in given instance $\mathcal{I}$, completing the proof of the theorem.
3 gives us that $\mathcal{A}^{*}$ is $\mathrm{IEF}1$ in $\widehat{\mathcal{I}}_{\ell_{k}}$, for each $k\in\mathbb{N}$. Using this observation (for any $k$) and Lemma 4, we obtain that $(A^{*}_{1}\setminus\{m+1\},\ldots,A^{*}_{n}\setminus\{m+1\})$ is $\mathrm{IEF}1$ in the instance $\mathcal{I}$ as well.
It remains to prove that $(A^{*}_{1}\setminus\{m+1\},\ldots,A^{*}_{n}\setminus\{m+1\})\in\mathrm{OPT}(\textbf{{LP}}(w^{*},\eta,\mathcal{I}))$. Towards this, consider the instance $\mathcal{I}^{\prime}$ obtained by including item $(m+1)$ into $\mathcal{I}$ and keeping the values, $v_{i}(t)$s, of all the other items unchanged. We first show that the tuple $(\mathcal{A}^{*},p^{*})$ satisfies the complementary slackness condition for the primal-dual linear programs $\textbf{{LP}}(w^{*},\eta,\mathcal{I}^{\prime})$ and $\textbf{{Dual-LP}}(w^{*},\eta,\mathcal{I}^{\prime})$. Together with the fact that $\mathcal{A}^{*}$ and $p^{*}$ are feasible for the linear programs $\textbf{{LP}}(w^{*},\eta,\mathcal{I}^{\prime})$ and $\textbf{{Dual-LP}}(w^{*},\eta,\mathcal{I}^{\prime})$, respectively, we obtain that $\mathcal{A}^{*}\in\mathrm{OPT}(\textbf{{LP}}(w^{*},\eta,\mathcal{I}^{\prime}))$.
To prove that complementary slackness holds, consider any agent $i\in[n]$ and item $t\in[m]$. If we have $t\notin A^{*}_{i}$, then complementary slackness holds trivially. Otherwise, if $t\in A^{*}_{i}$, then using the fact that $(\mathcal{A}^{*},p^{\ell_{k}})$ are an optimal primal-dual pair (3), we have $p^{\ell_{k}}_{t}=\big((w_{\ell_{k}})_{i}+\eta\big)\widehat{v}^{\ell_{k}}_{i}(t)$, where $\widehat{v}^{\ell_{k}}_{i}(t)$ is the value of item $t$ for agent $i$ in the instance $\widehat{\mathcal{I}}_{\ell_{k}}$. Taking the limit on both sides, we get
$p^{*}_{t}=\lim_{k\to\infty}p^{\ell_{k}}_{t}=\lim_{k\to\infty}\big((w_{\ell_{k}})_{i}+\eta\big)\widehat{v}^{\ell_{k}}_{i}(t)=(w^{*}_{i}+\eta)v_{i}(t).$
A similar argument shows that the complementary slackness holds for the item $(m+1)$ as well. Hence, $(\mathcal{A}^{*},p^{*})$ satisfies complementary slackness, implying that $\mathcal{A}^{*}\in\mathrm{OPT}(\textbf{{LP}}(w^{*},\eta,\mathcal{I}^{\prime}))$.
Finally, note that, for all agents $i\in[n]$ and all items $t\in[m]$, the valuations $v_{i}(t)$s are the same between $\mathcal{I}$ and $\mathcal{I}^{\prime}$. In addition, the optimality of $\mathcal{A}^{*}$ in $\mathcal{I}^{\prime}$ implies that each item $t\in[m]$ is assigned to an agent in $\operatorname*{arg\,max}_{j\in[n]}\ (w^{*}_{i}+\eta)v_{i}(t)$. This per-item guarantee implies that $(A^{*}_{1}\setminus\{m+1\},\ldots,A^{*}_{n}\setminus\{m+1\})\in\mathrm{OPT}(\textbf{{LP}}(w^{*},\eta,\mathcal{I}))$. Hence, $(A^{*}_{1}\setminus\{m+1\},\ldots,A^{*}_{n}\setminus\{m+1\})$ is $\mathrm{fPO}$ in $\mathcal{I}$.
Overall, we have that $(A^{*}_{1}\setminus\{m+1\},\ldots,A^{*}_{n}\setminus\{m+1\})$ is both $\mathrm{IEF}1$ and $\mathrm{fPO}$ in the given instance $\mathcal{I}$. The theorem stands proved. ∎
8 $\mathrm{EF}$ and $\mathrm{fPO}$ Allocations for Divisible Mixed Manna
This section shows that our result for indivisible mixed manna (specifically, Theorem 18) implies the envy-free ($\mathrm{EF}$) and $\mathrm{fPO}$ guarantee for divisible mixed manna under additive valuations. As mentioned previously, this result for the divisible case provides an alternate proof of the existential guarantee from [Bogo2017].
To prove the fairness and efficiency guarantee for divisible mixed manna, we will use the following lemma, which shows that the set of $\mathrm{fPO}$ allocations form a closed set. [^14] Although this lemma seems standard, we could not find a reference, and hence, provide a proof here for completeness.
Lemma 19.
[…]
Proof.
Write $\mathcal{F}\subset\mathbb{R}^{nm}$ to denote the set of all fractional allocations in the given instance $\mathcal{I}$. Note that $\mathcal{F}$ is a polytope and has finitely many faces.
Moreover, any fractional allocation $\mathcal{X}\in\mathcal{F}$ is $\mathrm{fPO}$ iff there exists a weight vector $\lambda\in\mathbb{R}^{n}_{>0}$ (i.e., a vector with $\lambda_{i}>0$ for each $i$) such that $\mathcal{X}$ maximizes the weighted social welfare with respect to $\lambda$; see [SS22Sharing, Lemma 2.3] and [mas1995microeconomic, Proposition 16.E.2]. That is, $\mathcal{X}\in\operatorname*{arg\,max}_{\mathcal{Y}\in\mathcal{F}}\ \sum_{t=1}^{m}\sum_{i=1}^{n}\lambda_{i}v_{i}(t)Y_{i,t}$.
Since the weighted welfare maximization problem is a linear program, for any weight vector $\lambda\in\mathbb{R}^{n}_{>0}$, the set of welfare-maximizing allocations is a face of the polytope $\mathcal{F}$. Write $\mathcal{F}_{\lambda}$ to denote this face, $\mathcal{F}_{\lambda}=\operatorname*{arg\,max}_{\mathcal{Y}\in\mathcal{F}}\ \sum_{t=1}^{m}\sum_{i=1}^{n}\lambda_{i}v_{i}(t)Y_{i,t}$.
The above-mentioned characterization of $\mathrm{fPO}$ implies that the set all $\mathrm{fPO}$ fractional allocations $\mathscr{F}=\bigcup\limits_{\lambda\in\mathbb{R}^{n}_{>0}}\ \mathcal{F}_{\lambda}$. Note that this union is composed of finitely many faces; the polytope $\mathcal{F}$ itself has finite number of faces. Since every face of $\mathcal{F}$ is closed, we get that $\mathscr{F}$ is a finite union of closed sets, and hence, is closed as well. The lemma stands proved. ∎
The following theorem is the main result of this section. We employ our result for indivisible items, along with limit arguments, to establish this theorem for divisible mixed manna.
Theorem 20.
Every fair division instance $\mathcal{D}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$ with divisible mixed manna and additive valuations admits an $\mathrm{EF}$ and $\mathrm{fPO}$ allocation $\mathcal{X}^{*}\in[0,1]^{n\times m}$.
Proof.
Consider the following infinite sequence of instances with indivisible items: For each $k\in\mathbb{N}$, write $\mathcal{I}_{k}=\langle[n],[mk],\{v^{k}_{i}\}_{i=1}^{n}\rangle$ to denote an instance with $mk$ indivisible items and agents’ additive valuations $v^{k}_{i}(\cdot)$. For each (divisible) item $t\in[m]$ in the given instance $\mathcal{D}$, we include $k$ indivisible items in $\mathcal{I}_{k}$. In particular, for each $t\in[m]$, the $k$ indivisible items $(t-1)k+1,(t-1)k+2,\ldots,tk$ in $\mathcal{I}_{k}$ are obtained by dividing the $t$th item in $\mathcal{D}$ equally. Here, for each agent $i\in[n]$, in instance $\mathcal{I}_{k}$ the values $v^{k}_{i}((t-1)k+1)=v^{k}_{i}((t-1)k+2)=\ldots=v^{k}_{i}(tk)=v_{i}(t)/k$.
The following claim establishes useful connections between the instances $\mathcal{I}_{k}$ and $\mathcal{D}$.
Claim 4.
Let $\mathcal{F}\coloneqq\{\mathcal{X}\in[0,1]^{n\times m}\mid\sum_{i=1}^{n}X_{i,t}=1\text{ for all }t\in[m]\}$ be the set of fractional allocations in the given instance $\mathcal{D}$ and $\mathcal{F}_{k}=\{\mathcal{X}\in[0,1]^{n\times mk}\mid\sum_{i=1}^{n}X_{i,t}=1\text{ for all }t\in[mk]\}$ be the set of fractional allocations in instance $\mathcal{I}_{k}$. Then, there exists a surjective (onto) map $g:\mathcal{F}_{k}\mapsto\mathcal{F}$ with the property that, for each fractional allocation $\mathcal{Y}\in\mathcal{F}_{k}$, every agent receives the same value in $g(\mathcal{Y})\in\mathcal{F}$ as in $\mathcal{Y}$, i.e., $v_{i}\left(g(\mathcal{Y})_{i}\right)=v^{k}_{i}\left(Y_{i}\right)$ for each $i\in[n]$.
Proof.
We define $g:\mathcal{F}_{k}\mapsto\mathcal{F}$ as follows: for any fractional allocation $\mathcal{Y}\in\mathcal{F}_{k}$, set $g(\mathcal{Y})_{i,t}=\frac{1}{k}\sum_{\ell=1}^{k}Y_{i,(t-1)k+\ell}$, for each $i\in[n]$ and $t\in[m]$.
First, to establish that $g$ is surjective, fix any $\mathcal{Z}\in\mathcal{F}$. We will show that there exists an allocation $\mathcal{X}\in\mathcal{F}_{k}$ such that $\mathcal{Z}=g(\mathcal{X})$. Specifically, set $X_{i,(t-1)k+\ell}\coloneqq Z_{i,t}$, for each $i\in[n]$, $t\in[m]$, and $\ell\in[k]$. Indeed, with this definition, we have $\mathcal{X}\in\mathcal{F}_{k}$. Furthermore, $g(\mathcal{X})_{i,t}=\frac{1}{k}\sum_{\ell=1}^{k}X_{i,(t-1)k+\ell}=\frac{1}{k}\sum_{\ell=1}^{k}Z_{i,t}=Z_{i,t}$ for each $i\in[n]$ and $t\in[m]$. Hence, we have $\mathcal{Z}=g(\mathcal{X})$, i.e., $g$ is surjective.
We next complete the proof of the claim by proving that $g$ preserves agents’ values. For each $\mathcal{Y}\in\mathcal{F}_{k}$ and agent $i\in[n]$ it holds that
$v_{i}(g(\mathcal{Y})_{i})$ $=\sum_{t=1}^{m}v_{i}(t)\ g(\mathcal{Y})_{i,t}=\sum_{t=1}^{m}\sum_{\ell=1}^{k}\frac{v_{i}(t)}{k}Y_{i,(t-1)k+\ell}$
$=\sum_{j=1}^{n}\sum_{\ell=1}^{k}v^{k}_{i}((t-1)k+\ell)\ Y_{i,(t-1)k+\ell}$ ($v^{k}_{i}((t-1)k+\ell)=\frac{v_{i}(t)}{k}$)
$=v_{i}^{k}(Y_{i}).$
The claim stands proved. ∎
For each instance $\mathcal{I}_{k}$ in the infinite sequence $\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3},\ldots$, we can apply Theorem 18 to infer the existence of an allocation $\mathcal{A}^{k}$ that is $\mathrm{fPO}$ and $\mathrm{IEF}1$ in $\mathcal{I}_{k}$. For each $k\in\mathbb{N}$, write $\mathcal{Q}^{k}\in[0,1]^{n\times mk}$ to denote the fractional allocation representation of $\mathcal{A}^{k}$, i.e., $Q^{k}_{i,t}=1$ if $t\in A^{k}_{i}$ and, otherwise, $Q^{k}_{i,t}=0$. From the sequence of allocations $\left\{\mathcal{A}^{k}\right\}_{k\in\mathbb{N}}$—equivalently, $\left\{\mathcal{Q}^{k}\right\}_{k\in\mathbb{N}}$ —we construct the sequence of fractional allocations $\left\{\mathcal{X}^{k}\right\}_{k\in\mathbb{N}}$ in the given instance $\mathcal{D}$ by setting $\mathcal{X}^{k}\coloneqq g(\mathcal{Q}^{k})$ for each $k$; here $g$ is the map obtained in 4.
Since the set of fractional allocations in $\mathcal{D}$ is closed and bounded, the Bolzano-Weierstrass theorem implies that the sequence $\left\{\mathcal{X}^{k}\right\}_{k\in\mathbb{N}}$ has a convergent subsequence $\mathcal{X}^{a_{1}},\mathcal{X}^{a_{2}},\mathcal{X}^{a_{3}},\ldots$. Let $\mathcal{X}^{*}$ be the limit point of this subsequence.
We will show that the fractional allocation $\mathcal{X}^{*}\in[0,1]^{n\times m}$ is $\mathrm{EF}$ and $\mathrm{fPO}$ in the instance $\mathcal{D}$. The existence of such a fair and efficient allocation establishes the theorem.
We first prove that $\mathcal{X}^{*}$ is $\mathrm{EF}$ in $\mathcal{D}$. Here, we upper bound the maximum envy in the fractional allocations $\mathcal{X}^{a_{k}}\in[0,1]^{n\times m}$ that form the converging subsequence. Recall that $\mathcal{X}^{a_{k}}$ is obtained by applying map $g(\cdot)$ to the (fractional representation of) the allocation $\mathcal{A}^{a_{k}}$ that is $\mathrm{IEF}1$ allocation in the instance $\mathcal{I}_{a_{k}}$ (Theorem 18). Hence, in instance $\mathcal{I}_{a_{k}}$, for each agent $i\in[n]$, there exists a subset $S\subseteq[mk]$ of size at most one ($|S|\leq 1$) such that $v_{i}^{a_{k}}(A^{a_{k}}_{i}\triangle S)\geq\max_{j\in[n]}v_{i}^{a_{k}}(A^{a_{k}}_{j})$. That is, $v_{i}^{a_{k}}(A^{a_{k}}_{i})+\max_{s\in[mk]}\ |v_{i}^{a_{k}}(s)|\geq\max_{j\in[n]}v_{i}^{a_{k}}(A^{a_{k}}_{j})$. Therefore, the maximum envy under allocation $\mathcal{A}^{a_{k}}$ in instance $\mathcal{I}_{a_{k}}$ is upper bounded as follows
$\max_{i,j\in[n]}\left(v_{i}^{a_{k}}(A^{a_{k}}_{j})-v_{i}^{a_{k}}(A^{a_{k}}_{i})\right)\leq\max_{i\in[n],\ s\in[mk]}|v_{i}^{a_{k}}(s)|=\max_{i\in[n],\ t\in[m]}\frac{1}{a_{k}}|v_{i}(t)|\leq\max_{i\in[n],\ t\in[m]}\frac{1}{k}|v_{i}(t)|$ (29)
Fractional allocation $\mathcal{X}^{a_{k}}$ is obtained by applying map $g(\cdot)$ to $\mathcal{A}^{a_{k}}$. Hence, using 4, we extend inequality (29) to upper bound the maximum envy in $\mathcal{X}^{a_{k}}$:
$\max_{i,j\in[n]}\left(v_{i}(X^{a_{k}}_{j})-v_{i}(X^{a_{k}}_{i})\right)\leq\frac{1}{k}\left(\max_{i\in[n],\ t\in[m]}|v_{i}(t)|\right)$ (30)
Note that the maximum envy is a continuous function of the fractional allocations $\mathcal{X}\in[0,1]^{n\times m}$. In addition, the right-hand-side of equation (30) tends to zero as $k\to\infty$. Hence, via the Order Limit Theorem, we obtain that the maximum envy in $\mathcal{X}^{*}$ is at most zero: $\max_{i,j\in[n]}\left(v_{i}(X^{*}_{j})-v_{i}(X^{*}_{i})\right)\leq 0$. Therefore, as stated, $\mathcal{X}^{*}$ is $\mathrm{EF}$ in the given instance $\mathcal{D}$.
To prove that $\mathcal{X}^{*}$ is $\mathrm{fPO}$ in the instance $\mathcal{D}$, we will show that all the fractional allocations $\mathcal{X}^{a_{k}}\in[0,1]^{n\times m}$ in the converging subsequence are $\mathrm{fPO}$ in $\mathcal{D}$. Since the set of $\mathrm{fPO}$ allocations in $\mathcal{D}$ is a closed set (Lemma 19), we get that the limit point of the subsequence, $\mathcal{X}^{*}$, is also $\mathrm{fPO}$.
Recall that, for each $k\in\mathbb{N}$, the fractional allocation $\mathcal{X}^{a_{k}}$ is obtained by applying the map $g(\cdot)$ to the (fractional representation of) the allocation $\mathcal{A}^{a_{k}}$ that is $\mathrm{fPO}$ allocation in the instance $\mathcal{I}_{a_{k}}$ (Theorem 18). Now, consider any fractional allocation $\mathcal{Z}$ in $\mathcal{D}$ – we will show that $\mathcal{Z}$ cannot Pareto dominate $\mathcal{X}^{a_{k}}$. Given that the map $g$ is surjective (4), there exists a fractional allocation $\mathcal{Y}$ in $\mathcal{I}_{a_{k}}$ such that $\mathcal{Z}=g(\mathcal{Y})$. Since $g$ preserves the valuations of all the $n$ agents (4) and $\mathcal{Y}$ does not Pareto dominate $\mathcal{I}_{a_{k}}$, we have that $\mathcal{Z}$ does not Pareto dominate $\mathcal{X}^{a_{k}}$. Hence, the fractional allocations $\mathcal{X}^{a_{k}}$, along with the limit $\mathcal{X}^{*}$, are $\mathrm{fPO}$.
Overall, we obtain the existence of an $\mathrm{EF}$ and $\mathrm{fPO}$ allocation of the divisible mixed manna in the given instance $\mathcal{D}$. This completes the proof of the theorem. ∎
9 Conclusion and Future Work
The current work nearly resolves, in the positive, the open problem concerning the existence of $\mathrm{EF}1$ and $\mathrm{PO}$ allocations for indivisible mixed manna under additive valuations. In addition to the $\mathrm{IEF}1$ and $\mathrm{PO}$ result for mixed manna, we provide a unified framework for establishing the $\mathrm{EF}1$ and $\mathrm{PO}$ guarantee for indivisible goods and indivisible chores, respectively. Our techniques also encapsulate the divisible-items setting – we develop an alternate proof of the established $\mathrm{EF}$ and $\mathrm{PO}$ guarantee for divisible mixed manna.
Resolving the fairness and efficiency question exactly with the $\mathrm{EF}1$ criterion remains an interesting direction for future work. We note that this would require strengthening the underlying KKM instantiation itself. In particular, via examples, one can show that starting from Theorem 10 the best one can hope for is an $\mathrm{IEF}1$ (or $\mathrm{XEF}1$) guarantee. Developing pseudo-polynomial time algorithms for finding fair and efficient allocation of indivisible mixed manna (or even indivisible chores) is another interesting direction.
Appendix A $\mathrm{EF}1$ and $\mathrm{PO}$ for Indivisible Goods
This section provides a sketch of how the proof of Theorem 1 can be modified to establish the existence of $\mathrm{PO}$ and $\mathrm{EF}1$ allocations of indivisible goods. In particular, we will show the existence of mixed-manna allocations that are $\mathrm{PO}$, and extrospectively envy-free up to one item ($\mathrm{XEF}1$); this fairness notion is defined next.
Definition 4.
In an instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$, an allocation $\mathcal{A}=(A_{1},\ldots,A_{n})$ is said to be extrospectively envy-free up to one item ($\mathrm{XEF}1$) if, for all agents $i,j\in[n]$, there exists a subset $S\subseteq[m]$ of size at most one ($|S|\leq 1$) such that $v_{i}(A_{i})\geq v_{i}(A_{j}\triangle S)$.
To avoid repeating the arguments used in the proof of Theorem 1, we will primarily focus on the modifications required to establish the existence of $\mathrm{PO}$ and $\mathrm{XEF}1$ allocations.
To prove this existential result, some changes are required in the construction of the non-degenerate instance $\widehat{\mathcal{I}}=\langle[n],[\widehat{m}],\{\widehat{v}_{i}\}_{i=1}^{m}\rangle$ (see Section 4.1 for comparison) and in the primal and dual programs (see Section 4.2). We begin by describing the changes in $\widehat{\mathcal{I}}$; with a slight abuse of notations we continue to denote the non-degenerate instance by $\widehat{\mathcal{I}}$. We set $\widehat{m}\coloneqq m+n$, i.e., we add $n$ auxiliary items to the given instance $\mathcal{I}$. For each $i\in[n]$, the valuations of the new item, $(m+i)$, is set as follows: $\widehat{v}_{i}(m+i)=\lambda/2$ and $\widehat{v}_{j}(m+i)=0$, for all $j\in[n]\setminus\{i\}$. For the given items $t\in[m]$ and any agent $i\in[n]$, if $v_{i}(t)=0$, we set $\widehat{v}_{i}(t)=0$, i.e., leave the zero values unchanged. Otherwise, if $v_{i}(t)\neq 0$, then we set $\widehat{v}_{i}(t)=v_{i}(t)-\overline{\varepsilon}_{i,t}$. The perturbations $\overline{\varepsilon}_{i,t}$ are drawn independently from the uniform distribution $\mathrm{Uniform}[0,\overline{\varepsilon}]$, where $\overline{\varepsilon}>0$ is fixed to satisfy
$\overline{\varepsilon}<\frac{\lambda\omega}{16mn^{2}(m+n)\ \max\limits_{i\in[n],t\in[m]}|v_{i}(t)|}$ (31)
Here, $\lambda$ and $\omega$ are defined as in Section 4.1. We will now describe the modified linear programs, denoted by $\overline{\textbf{{LP}}}$ and $\overline{\textbf{{Dual-LP}}}$. Their construction will utilize the following constant
$\overline{\eta}\coloneqq\frac{\lambda}{2n(m+n)\max\limits_{i\in[n],\ t\in[\widehat{m}]}|\widehat{v}_{i}(t)|}$ (32)
The constants $\overline{\varepsilon}$ and $\overline{\eta}$ satisfy the following analogue of Equation 5
$\overline{\varepsilon}<\frac{\overline{\eta}\omega}{4mn}$ (33)
[…]
The following lemma is a counterpart of Lemma 4; its proof follows similar arguments and is omitted.
Lemma 21.
If an allocation $\mathcal{A}=(A_{1},\ldots,A_{n})$ is $\mathrm{XEF}1$ in the non-degenerate instance $\widehat{\mathcal{I}}=\langle[n],[\widehat{m}],\{\widehat{v}_{i}\}_{i=1}^{n}\rangle$, then the allocation $(A_{1}\setminus\{(m+i)\}_{i=1}^{n},A_{2}\setminus\{(m+i)\}_{i=1}^{n},\ldots,A_{n}\setminus\{(m+i)\}_{i=1}^{n})$ is $\mathrm{XEF}1$ in the given instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$.
We will now state an analogue of Lemma 5. While the proof here follows an approach similar to that of Lemma 5, key technical aspects differ and, hence, we include the proof below.
Lemma 22.
If an allocation $\mathcal{A}=(A_{1},\ldots,A_{n})\in\mathrm{OPT}(\overline{\textbf{{LP}}}(w,\overline{\eta}))$, for any $w\in\Delta_{n-1}$, then the allocation $(A_{1}\setminus\{(m+i)\}_{i=1}^{n},A_{2}\setminus\{(m+i)\}_{i=1}^{n},\ldots,A_{n}\setminus\{(m+i)\}_{i=1}^{n})$ is $\mathrm{PO}$ in the given instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$.
Proof.
Define an intermediate instance $\mathcal{I}^{\prime}=\langle[n],[m+n],\{v^{\prime}_{i}\}_{i=1}^{n}\rangle$ wherein $v^{\prime}_{i}(t)=v_{i}(t)$, for all agents $i\in[n]$ and given items $t\in[m]$. In addition, set $v^{\prime}_{i}(m+t)=\widehat{v}_{i}(m+t)$ for each $i,t\in[n]$. Note that instance $\mathcal{I}^{\prime}$ is constructed by adding the auxiliary items $m+1,\ldots,m+n$ to $\mathcal{I}$ and keeping the valuations of the original $m$ items unchanged. Instance $\mathcal{I}^{\prime}$ is defined for the purposes of analysis: We will first show that $\mathcal{A}$ is $\mathrm{PO}$ in $\mathcal{I}^{\prime}$. Then, we will conclude the proof by showing that $(A_{1}\setminus\{(m+i)\}_{i=1}^{n},\ldots,A_{n}\setminus\{(m+i)\}_{i=1}^{n})$ is $\mathrm{PO}$ in the given instance $\mathcal{I}$.
Towards a contradiction, we assume that there exists an allocation $\mathcal{B}=(B_{1},\ldots,B_{n})$ that Pareto dominates $\mathcal{A}$ in $\mathcal{I}^{\prime}$. That is, $v^{\prime}_{i}(B_{i})\geq v^{\prime}_{i}(A_{i})$, for all $i\in[n]$, and $v^{\prime}_{a}(B_{a})>v^{\prime}_{a}(A_{a})$ for some $a\in[n]$. Without loss of generality, we also assume that $\mathcal{B}$ is Pareto efficient in $\mathcal{I}^{\prime}$; this ensures that, for each $i\in[n]$, the auxiliary item $(m+i)$ is allocated to agent $i$ in $\mathcal{B}$. Additionally, note that each auxiliary item $(m+i)$ also gets allocated to agent $i$ in $\mathcal{A}$ since $\mathcal{A}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(w,\overline{\eta}))$. These observations imply that
$0$ $<\sum_{i=1}^{n}v^{\prime}_{i}(B_{i})-\sum_{i=1}^{n}v^{\prime}_{i}(A_{i})$
$=n\frac{\lambda}{2}+\sum_{i=1}^{n}v^{\prime}_{i}(B_{i}\setminus\{(m+i)\}_{i=1}^{n})-n\frac{\lambda}{2}-\sum_{i=1}^{n}v^{\prime}_{i}(A_{i}\setminus\{(m+i)\}_{i=1}^{n})$ ($v^{\prime}_{i}(m+i)=\lambda/2$ for all $i$)
$=\sum_{i=1}^{n}v_{i}(B_{i}\setminus\{(m+i)\}_{i=1}^{n})-\sum_{i=1}^{n}v_{i}(A_{i}\setminus\{(m+i)\}_{i=1}^{n})$ (34)
That is, $\sum_{i=1}^{n}v_{i}(B_{i}\setminus\{(m+i)\}_{i=1}^{n})-\sum_{i=1}^{n}v_{i}(A_{i}\setminus\{(m+i)\}_{i=1}^{n})>0$. Using this strict inequality and the definition of $\omega$, we obtain
$\omega\leq\sum_{i=1}^{n}v_{i}(B_{i}\setminus\{m+i\}_{i=1}^{n})-\sum_{i=1}^{n}v_{i}(A_{i}\setminus\{m+i\}_{i=1}^{n})=\sum_{i=1}^{n}v^{\prime}_{i}(B_{i})-\sum_{i=1}^{n}v^{\prime}_{i}(A_{i})$
Next, we use the welfare optimality of $\mathcal{A}$ in $\widehat{\mathcal{I}}$ (and under the weight vector $w$) to arrive at a contradiction. In particular, $\mathcal{A}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(w,\overline{\eta}))$ implies that $\sum_{i=1}^{n}(w_{i}+\overline{\eta})^{-1}\widehat{v}_{i}(B_{i})\leq\sum_{i=1}^{n}(w_{i}+\overline{\eta})^{-1}\widehat{v}_{i}(A_{i})$. For each subset $S\subseteq[\widehat{m}$] and each agent $i$, it holds that $v^{\prime}_{i}(S)-\overline{\varepsilon}|S|\leq\widehat{v}_{i}(S)\leq v^{\prime}_{i}(S)$. Therefore, we have $\sum_{i=1}^{n}(w_{i}+\overline{\eta})^{-1}(v^{\prime}_{i}(B_{i})-\overline{\varepsilon}|B_{i}|)\leq\sum_{i=1}^{n}(w_{i}+\overline{\eta})^{-1}v^{\prime}_{i}(A_{i})$. This inequality reduces to
$\sum_{i=1}^{n}\frac{1}{(w_{i}+\overline{\eta})}(v^{\prime}_{i}(B_{i})-v^{\prime}_{i}(A_{i}))$ $\leq\overline{\varepsilon}\sum_{i=1}^{n}\frac{1}{(w_{i}+\overline{\eta})}|B_{i}|$
$\leq\overline{\varepsilon}\sum_{i=1}^{n}\frac{1}{\overline{\eta}}|B_{i}|$ ($w_{i}\geq 0$)
$=\frac{\overline{\varepsilon}(m+n)}{\overline{\eta}}$ (36)
We can lower bound the left-hand-side of inequality (36) as follows
$\sum_{i=1}^{n}\frac{1}{(w_{i}+\overline{\eta})}(v^{\prime}_{i}(B_{i})-v^{\prime}_{i}(A_{i}))$ $\geq\frac{1}{2}\sum_{i=1}^{n}(v^{\prime}_{i}(B_{i})-v^{\prime}_{i}(A_{i}))$ ($v^{\prime}_{i}(B_{i})-v^{\prime}_{i}(A_{i})\geq 0$ and $w_{i}+\overline{\eta}\leq 2$)
$\geq\frac{\omega}{2}$ (via (A))
The last inequality and equation (36) imply $\frac{\omega}{2}\leq\frac{\overline{\varepsilon}(m+n)}{\overline{\eta}}$, or equivalently, $\frac{\overline{\eta}\omega}{2(m+n)}\leq\overline{\varepsilon}$. This bound, however, contradicts the strict inequality (33), which states that $\overline{\varepsilon}<\frac{\overline{\eta}\omega}{4mn}$. [^15] We have $n,m\geq 2$. Therefore, by way of contradiction, we obtain that no allocation $\mathcal{B}$ Pareto dominates $\mathcal{A}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(w,\overline{\eta}))$ in the instance $\mathcal{I}^{\prime}$, i.e., $\mathcal{A}$ is $\mathrm{PO}$ in $\mathcal{I}^{\prime}$.
Finally, we will use the Pareto optimality of $\mathcal{A}$ in $\mathcal{I}^{\prime}$ to show that the allocation $(A_{1}\setminus\{(m+i)\}_{i=1}^{n},\ldots,A_{n}\setminus\{(m+i)\}_{i=1}^{n})$ is $\mathrm{PO}$ in the given instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$. Assume, towards a contradiction, that allocation $(D_{1},\ldots,D_{n})$ Pareto dominates $(A_{1}\setminus\{(m+i)\}_{i=1}^{n},\ldots,A_{n}\setminus\{(m+i)\}_{i=1}^{n})$ in $\mathcal{I}$. As observed above, for each $i\in[n]$, the item $(m+i)$ is allocated to agent $i$ in allocation $\mathcal{A}$. Hence, our assumption would imply that $(D_{1}\cup\{(m+1)\},\ldots,D_{i}\cup\{(m+i)\},\ldots,D_{n}\cup\{(m+n)\})$ Pareto dominates $(A_{1},\ldots,A_{i},\ldots,A_{n})$ in $\mathcal{I}^{\prime}$. Since this contradicts the already established Pareto optimality of $\mathcal{A}$ in $\mathcal{I}^{\prime}$, we obtain that $(A_{1}\setminus\{(m+i)\}_{i=1}^{n},\ldots,A_{n}\setminus\{(m+i)\}_{i=1}^{n})$ is $\mathrm{PO}$ in the instance $\mathcal{I}$. The lemma stands proved. ∎
Another key change is how we apply the KKM theorem. Here, we define the sets $\overline{C}_{1},\overline{C}_{2},\ldots,\overline{C}_{n}\subseteq\Delta_{n-1}$ considering $\operatorname*{arg\,min}$ over the optimal prices (instead of $\operatorname*{arg\,max}$ as in Equation 10).
[…]
Note that each $\overline{C}_{i}$ is the set of weights $w\in\Delta_{n-1}$ for which there is an optimal allocation $\mathcal{A}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(w,\overline{\eta}))$ wherein $i$’s bundle has the lowest price, under the optimal $p$. This defining condition, that agent $i$ is worst off in terms of optimal bundle prices, is in complete contrast to the one imposed in Section 5 (Equation 10). Nonetheless, we show below that this somewhat counterintuitive condition leads to the desired $\mathrm{XEF}1$ guarantee.
The following lemma asserts that the sets $\overline{C}_{1},\ldots,\overline{C}_{n}\subseteq\Delta_{n-1}$ are closed; the proof of this lemma is similar to that of Lemma 8.
Lemma 23 (Closedness).
For each $i\in[n]$, the set $\overline{C}_{i}\subseteq\Delta_{n-1}$ is closed.
We will now prove that the sets $\overline{C}_{1},\ldots,\overline{C}_{n}$ satisfy the KKM covering condition; the proof of this lemma is notably different from its counterpart, Lemma 9.
Lemma 24 (KKM Covering).
For any $w\in\Delta_{n-1}$ there exists an $i\in\mathrm{supp}(w)$ such that $w\in\overline{C}_{i}$.
Proof.
Fix a weight vector $w\in\Delta_{n-1}$. We consider two cases based on whether $\mathrm{supp}(w)=[n]$ or $\mathrm{supp}(w)\subsetneq[n]$.
Case I: $\mathrm{supp}(w)=[n]$. Consider any allocation $\mathcal{A}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(w,\overline{\eta}))$ and optimal prices $p=\mathrm{OPT}(\overline{\textbf{{Dual-LP}}}(w,\overline{\eta}))$. Let $i^{\prime}\in\operatorname*{arg\,min}_{j\in[n]}p(A_{j})$. In the current case $\mathrm{supp}(w)=[n]$ and, hence, $i^{\prime}\in\mathrm{supp}(w)$. Additionally, we have $w\in\overline{C}_{i^{\prime}}$, since $i^{\prime}\in\operatorname*{arg\,min}_{j\in[n]}p(A_{j})$. Hence, for vectors $w$ with full support, the KKM covering condition holds.
Case II: $\mathrm{supp}(w)\subsetneq[n]$. In this case $n]\setminus\mathrm{supp}(w)\neq\emptyset$. Write optimal prices $p=\mathrm{OPT}(\overline{\textbf{{Dual-LP}}}(w,\overline{\eta}))$ and consider any allocation $\mathcal{A}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(w,\overline{\eta}))$. We will show that there exists an $i^{\prime}\in\mathrm{supp}(w)$ with the property that $i^{\prime}\in\operatorname*{arg\,min}_{j\in[n]}\ p(A_{j})$. For such an agent $i^{\prime}$, we have $w\in C_{i^{\prime}}$, and, hence, we will obtain KKM covering in the current case as well.
Define $\coloneqq\{i\in[n]\mid w_{i}\geq 1/n\}; HH\neq\emptyset$ since $\sum_{i=1}^{n}w_{i}=1$. Also, the set of agents $[n]$ partitions into three disjoint subsets: $n]\setminus\mathrm{supp}(w)$, $H$, and $\mathrm{supp}(w)\setminus H$.
Let $h\in\operatorname*{arg\,max}_{i\in H}w_{i}$ be an agent with the highest weight value; note that $h\in H\subseteq\mathrm{supp}(w)$.
To show that there exists an $i^{\prime}\in\mathrm{supp}(w)$ such that $w\in\overline{C}_{i^{\prime}}$, we will first show that $\min\limits_{i\in[n]\setminus\mathrm{supp}(w)}p(A_{i})\geq p(A_{h})$. This will conclude the proof, since we will obtain
$\min\limits_{i\in[n]\setminus\mathrm{supp}(w)}p(A_{i})\geq p(A_{h})\geq\min\limits_{j\in\mathrm{supp}(w)}p(A_{j}).$
Here, the final inequality follows from $h\in\mathrm{supp}(w)$. This will in turn imply the existence of an $i^{\prime}\in\mathrm{supp}(w)$ such that $i^{\prime}\in\operatorname*{arg\,min}_{j\in[n]}p(A_{j})$, i.e., $w\in\overline{C}_{i^{\prime}}$, completing the proof.
Hence, in the remainder of the proof, we will show that $\min\limits_{i\in[n]\setminus\mathrm{supp}(w)}p(A_{i})\geq p(A_{h})$. Towards this, we will prove the following two inequalities
$\min\limits_{i\in[n]\setminus\mathrm{supp}(w)}p(A_{i})\geq\lambda/2\overline{\eta}$ (38)
$\lambda/2\overline{\eta}\geq p(A_{h})$ (39)
Together inequalities (38) and (39) imply the desired bound $\min\limits_{i\in[n]\setminus\mathrm{supp}(w)}p(A_{i})\geq\lambda/2\overline{\eta}\geq p(A_{h})$.
[…]
We will now show that no chore $c\in C$ is allocated to the agent $i\in[n]\setminus\mathrm{supp}(w)$. To this end, it is suffices to show that $\widehat{v}_{h}(c)/(w_{h}+\overline{\eta})>\widehat{v}_{i}(c)/\overline{\eta}$, or equivalently $\overline{\eta}\widehat{v}_{h}(c)>(w_{h}+\overline{\eta})\widehat{v}_{i}(c)$. To establish this, note that
$(w_{h}+\overline{\eta})|\widehat{v}_{i}(c)|$ $>|\widehat{v}_{i}(c)|/n$ ($w_{h}\geq 1/n$ and $\overline{\eta}>0$)
$\geq\lambda/n$ ($|\widehat{v}_{i}(c)|=|v_{i}(c)-\varepsilon_{i,c}|\geq\lambda$)
$>\overline{\eta}\ \max_{a,b}|\widehat{v}_{a}(b)|$ (Equation 32)
$\geq\overline{\eta}\ |\widehat{v}_{h}(c)|.$
Since $\widehat{v}_{h}(c),\widehat{v}_{i}(c)<0$, we get the desired inequality $\overline{\eta}\widehat{v}_{h}(c)>(w_{h}+\overline{\eta})\widehat{v}_{i}(c)$.
Establishing Inequality (39). If an item $t\in[\widehat{m}$] is allocated to agent $h$, then we have $p_{t}=\widehat{v}_{h}(t)/(w_{h}+\overline{\eta})\leq\max_{a,b}|{\widehat{v}}_{a}(b)|/(w_{h}+\overline{\eta})<\max_{a,b}|{\widehat{v}}_{a}(b)|/w_{h}\leq n\max_{a,b}|{\widehat{v}}_{a}(b)|$, where the final inequality follows from $w_{h}\geq 1/n$. Therefore, $p(A_{h})\leq\sum_{t\in A_{h}}p_{t}\leq(m+n)\ n\max_{a,b}|{\widehat{v}}_{a}(b)|$. Finally, using the definition of $\overline{\eta}$ (Equation 32), we obtain the inequality (39): $p(A_{h})\leq n(m+n)\max_{a,b}|{\widehat{v}}_{a}(b)|=\lambda/2\overline{\eta}$.
This completes the proof of the lemma. ∎
Lemmas 23 and 24 together give us the following analogue of Theorem 10.
Theorem 25.
There exists a point $\overline{w}^{*}\in\Delta_{n-1}$ with the property that, for each $i\in[n]$, we have an allocation $\mathcal{A}^{(i)}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(\overline{w}^{*},\overline{\eta}))$ wherein $i\in\operatorname*{arg\,min}_{j\in[n]}p^{*}(A^{(i)}_{j})$; here $p^{*}=\mathrm{OPT}(\overline{\textbf{{Dual-LP}}}(\overline{w}^{*},\overline{\eta}))$.
The subsequent argument to prove the existence of $\mathrm{PO}$ and $\mathrm{XEF}1$ allocation also closely follows Section 6. Let $\overline{w}^{*}\in\Delta_{n-1}$ be the weight vector whose existence is guaranteed in Theorem 25, and let $p^{*}=\mathrm{OPT}(\overline{\textbf{{Dual-LP}}}(\overline{w}^{*},\overline{\eta}))$ be the associated optimal prices.
For each agent $i\in[n]$, write $\overline{L}_{i}$ to denote the set of items that must be allocated to agent $i$ in any allocation in $\mathrm{OPT}(\overline{\textbf{{LP}}}(\overline{w}^{*},\overline{\eta}))$, i.e., $\overline{L}_{i}\coloneqq\left\{s\in[\widehat{m}]\mid\frac{\widehat{v}_{i}(s)}{(\overline{w}^{*}_{i}+\overline{\eta})}>\frac{\widehat{v}_{j}(s)}{(\overline{w}^{*}_{j}+\overline{\eta})}\text{ for all }j\neq i\right\}$. Denote by $\overline{\mathcal{L}}=(\overline{L}_{1},\ldots,\overline{L}_{n})$ the partial allocation formed by these bundles and let $\overline{U}\coloneqq[\widehat{m}]\setminus\left(\cup_{j=1}^{n}\overline{L}_{j}\right)$.
Next, we consider a bipartite graph $\overline{H}=([n]\cup\overline{U},\overline{E})$, with $[n]$ and $\overline{U}$ as the left and right part, respectively. Edge $(i,t)\in\overline{E}\subseteq[n]\times\overline{U}$ is included in the bipartite graph if and only if $p^{*}_{t}=\frac{\widehat{v}_{i}(t)}{\overline{w}^{*}_{i}+\overline{\eta}}$. The vertices adjacent to any vertex $x$ in $\overline{H}$ will be denoted by $\overline{\Gamma}(x)$. Arguments similar to the ones used in Section 6 establish the following lemma for $\overline{H}$.
Lemma 26.
The bipartite graph $\overline{H}=([n]\cup U,\overline{E})$ is a forest.
Define collection of complete allocations $\overline{O}^{*}\coloneqq\left\{(B_{1},\ldots,B_{n})\in\Pi_{n}([\widehat{m}])\mid\overline{L}_{i}\subseteq B_{i}\subseteq\overline{L}_{i}\cup\overline{\Gamma}(i)\text{ for each }i\in[n]\right\}$. Analogous to Lemma 12, we have the following result for $\overline{O}^{*}$.
Proposition 4.
Any allocation $\mathcal{B}\in\overline{O}^{*}$ if and only if $\mathcal{B}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(\overline{w}^{*},\overline{\eta}))$.
Max-Min Allocations.
We now define some key constructs for the rest of the analysis. Note that while the high-level structure of the proof here is similar to Section 6.2, multiple constructs differ from the previous ones. In particular, the price threshold $\overline{\tau}$ (defined below) has a max-min formulation in contrast to $\tau$ defined Equation 19.
$\overline{\tau}\coloneqq\max_{(B_{1},\ldots,B_{n})\in\overline{O}^{*}}\ \ \min_{i\in[n]}p^{*}(B_{i}).$
Define set $\overline{\mathcal{M}}^{*}\coloneqq\{(B_{1},\ldots,B_{n})\in\overline{O}^{*}\mid\min_{j\in[n]}p^{*}(B_{j})=\overline{\tau}\}$ as the collection of all max-min allocations. Additionally, for each agent $i\in[n]$ and subsets $S$, we define
$p^{-}_{i}(S)\coloneqq\min\limits_{\begin{subarray}{c}T\subseteq\Gamma(i),\\ |T|\leq 1\end{subarray}}p^{*}(S\triangle T).$
We also define, for each allocation $\mathcal{B}=(B_{1},\ldots,B_{n})$, the potential $\varphi^{-}(\mathcal{B})$, which counts the number of agents whose bundle prices can be toggled to $\overline{\tau}$ or below.
$\varphi^{-}(\mathcal{B})\coloneqq\left|\{i\in[n]\mid p_{i}^{-}(B_{i})\leq\overline{\tau}\}\right|.$
The following theorem is a key result of this section, and it shows that there exists a max-min allocation $\mathcal{A}^{*}$ wherein the bundle prices of all the agents can be toggled to $\overline{\tau}$ or below.
Theorem 27.
There exists an allocation $\mathcal{A}^{*}\in\overline{\mathcal{M}}^{*}$ such that $\varphi^{-}(\mathcal{A}^{*})=n$.
Before proving Theorem 27 (in Section A.1 below), we use this result to establish the main result of this section, the existence of $\mathrm{XEF}1$ and $\mathrm{PO}$ allocations.
Theorem 28.
Every fair division instance $\mathcal{I}=\langle[n],[m],\{v_{i}\}_{i=1}^{n}\rangle$, with indivisible mixed manna and additive valuations, admits an allocation that is both extrospectively envy-free up to one item ($\mathrm{XEF}1$) and Pareto efficient ($\mathrm{PO}$).
Proof.
Consider the weight vector $\overline{w}^{*}\in\Delta_{n-1}$ whose existence is implied by Theorem 25 and let $p^{*}=\allowbreak\mathrm{OPT}(\overline{\textbf{{Dual-LP}}}(\overline{w}^{*},\overline{\eta}))$. Let $\mathcal{A}^{*}\in\overline{\mathcal{M}}^{*}$ be the allocation whose existence is implied by Theorem 27.
First, note that $\mathcal{A}^{*}\in\overline{\mathcal{M}}^{*}\subseteq\overline{O}^{*}$, therefore by Proposition 4, we get $\mathcal{A}^{*}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(\overline{w}^{*},\overline{\eta}))$.
Next, we will show that $\mathcal{A}^{*}$ is $\mathrm{XEF}1$ for the instance $\overline{\mathcal{I}}$. By Theorem 27, we have $p^{-}_{j}(A^{*}_{j})\leq\overline{\tau}$ for all $j\in[n]$. Since, $\mathcal{A}^{*}\in\overline{\mathcal{M}}^{*}$ we also have $\overline{\tau}=\min_{i\in[n]}\ p^{*}(A^{*}_{i})$. Hence, for any pair of agents $i,j\in[n]$, we have
$\min\limits_{\begin{subarray}{c}T\subseteq\Gamma(j),\\ |T|\leq 1\end{subarray}}p^{*}(A^{*}_{j}\triangle T)=p^{-}_{j}(A^{*}_{j})\leq\overline{\tau}\leq p^{*}(A^{*}_{i}).$ (40)
Recall that $\mathcal{A}^{*}\in\overline{O}^{*}$ and, hence, $\mathcal{A}^{*}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(\overline{w}^{*},\overline{\eta}))$ (Proposition 4). Therefore, for each item $t\in A^{*}_{i}\subseteq\overline{L}_{i}\cup\overline{\Gamma}(i)$ it holds that $p^{*}_{t}=\widehat{v}_{i}(t)/(\overline{w}^{*}_{i}+\overline{\eta})$. Furthermore, for each item $s\in A^{*}_{j}\Delta T$, we have $p^{*}_{s}=\max_{a\in[n]}\widehat{v}_{a}(s)/(\overline{w}^{*}_{a}+\overline{\eta})\geq\widehat{v}_{i}(s)/(\overline{w}^{*}_{i}+\overline{\eta})$. Using these bounds with Equation 40, we obtain
$\frac{1}{\overline{w}^{*}_{i}+\overline{\eta}}\ \min\limits_{\begin{subarray}{c}T\subseteq\Gamma(j),\\ |T|\leq 1\end{subarray}}\widehat{v}_{i}(A^{*}_{j}\triangle T)\leq\min\limits_{\begin{subarray}{c}T\subseteq\Gamma(j),\\ |T|\leq 1\end{subarray}}p^{*}(A^{*}_{j}\triangle T)\leq p^{*}(A^{*}_{i})=\frac{1}{\overline{w}^{*}_{i}+\overline{\eta}}\ \widehat{v}_{i}(A^{*}_{i}).$ (41)
Since $(\overline{w}^{*}_{i}+\overline{\eta})>0$, we get that $\mathcal{A}^{*}$ is $\mathrm{XEF}1$ in the instance $\widehat{\mathcal{I}}$: For each pair of agents $i,j\in[n]$, there exists a a subset $T\subseteq\overline{\Gamma}(j)\subseteq[\widehat{m}$] of size at most one ($|T|\leq 1$) such that $\widehat{v}_{i}(A^{*}_{i})\geq\widehat{v}_{i}(A^{*}_{j}\triangle T)$.
Therefore, we have that the allocation $\mathcal{A}^{*}$ is $\mathrm{XEF}1$ in $\overline{\mathcal{I}}$ and $\mathcal{A}^{*}\in\mathrm{OPT}(\overline{\textbf{{LP}}}(\overline{w}^{*},\overline{\eta}))$. Finally, Lemmas 21 and 22 imply that the allocation $(A^{*}_{1}\setminus\{m+1\},A^{*}_{2}\setminus\{m+1\},\ldots,A^{*}_{n}\setminus\{m+1\})$ is $\mathrm{XEF}1$ and $\mathrm{PO}$ in the given instance $\mathcal{I}$. This completes the proof of the theorem. ∎
A.1 Proof of Theorem 27
1
2
3
Input : Allocation $\mathcal{S}=(S_{1},\dots,S_{n})\in\overline{\mathcal{M}}^{*}$ (i.e., $\min_{i}p^{*}(S_{i})=\overline{\tau}$) with $\varphi^{-}(\mathcal{S})<n$.
Output : Allocation $\widehat{\mathcal{S}}=(\widehat{S}_{1},\dots,\widehat{S}_{n})\in\overline{\mathcal{M}}^{*}$ satisfying $\varphi^{-}(\widehat{\mathcal{S}})>\varphi^{-}(\mathcal{S})$.
4
5Let $r\in[n]$ to be any agent with $p^{-}_{r}(S_{r})>\overline{\tau}$. Initialize set $\mathcal{Q}=\{r\}$.
6
7Write $R$ to denote the tree in bipartite graph (forest) $\overline{H}$ rooted at $r$. For each node $x$ in the tree $R$, let $\mathrm{par}(x)$ denote the parent of $x$. In addition, $\mathrm{ch}(x)=\Gamma(x)\setminus\{\mathrm{par}(x)\}$ are the children of $x$ in the tree. Also, for each agent $i$ in the tree, $\overline{\mathrm{ch}}(i)=\cup_{t\in\mathrm{ch}(i)}\ \mathrm{ch}(t)$ denotes the agents that are immediate descendants of $i$.
8 Initialize $\widehat{S}_{j}=S_{j}$ for all agents $j\in[n]$.
9
101exwhile $\mathcal{Q}\neq\emptyset$ do
11   Let $i\in\mathcal{Q}$ be any agent in $\mathcal{Q}$. Update $\mathcal{Q}\leftarrow\mathcal{Q}\setminus\{i\}$.
12
13  1ex
14  Set $X\subseteq\mathrm{ch}(i)$ to be an item subset that satisfies: (i) $p^{-}_{i}(\widehat{S}_{i}\triangle X)\leq\overline{\tau}$, (ii) $p^{*}(\widehat{S}_{i}\triangle X)\geq\overline{\tau}$, and (iii) $p^{*}(\widehat{S}_{i}\triangle\{t\})<p^{*}(\widehat{S}_{i})$ for each $t\in X$. {Existence of $X$ is guaranteed via Lemma 29.}
15
16  1exfor all items $t\in X$ do
17      If item $t\in\widehat{S}_{i}$, then set $a$ to be any agent in $\mathrm{ch}(t)\neq\emptyset$. Otherwise, if $t\notin\widehat{S}_{i}$, then let $a\in\mathrm{ch}(t)$ be the agent such that $t\in\widehat{S}_{a}$. {Such an agent $a\in\overline{\mathrm{ch}}(i)$ always exists.}
18      Update $\widehat{S}_{a}\leftarrow\widehat{S}_{a}\triangle\{t\}$.
19
20     1exIf $p^{-}_{a}(\widehat{S}_{a})>\overline{\tau}$, then include $a$ for processing: $\mathcal{Q}\leftarrow\mathcal{Q}\cup\{a\}$. {Otherwise, if $p^{-}_{a}(\widehat{S}_{a})\leq\overline{\tau}$, then we do not include $a$ in $\mathcal{Q}$.}
21
22
23  1exUpdate $\widehat{S}_{i}\leftarrow\widehat{S}_{i}\triangle X$
24
25return Allocation $\widehat{\mathcal{S}}=(\widehat{S}_{1},\ldots,\widehat{S}_{n})$.
Algorithm 2 $\mathrm{XEF}1$AugmentTree Algorithm
The proofs of the following three lemmas (Lemmas 29 to 31) follow along the lines of the proofs of Lemmas 14 to 16, and hence are omitted.
Lemma 29.
For any agent $i\in[n]$ in the rooted tree, let $S$ be an item subset such that $\overline{L}_{i}\subseteq S\subseteq\overline{L}_{i}\cup\overline{\Gamma}(i)$ and $p^{-}_{i}(S)>\overline{\tau}$. Then, there exists subset $X\subseteq\mathrm{ch}(i)$ satisfying the following properties
  1. (i)
    $p^{-}_{i}(S\triangle X)\leq\overline{\tau}$,
  2. (ii)
    $p^{*}(S\triangle X)\geq\overline{\tau}$, and
  3. (iii)
    $p^{*}(S\triangle\{t\})<p^{*}(S)$ for each $t\in X$.
Lemma 30.
Each iteration of the while-loop in the $\mathrm{XEF}1$AugmentTree Algorithm executes successfully and maintains an allocation $\widehat{\mathcal{S}}\in\overline{\mathcal{M}}^{*}$. Furthermore, at the end of the iteration (Line 2), for the agent $i$ selected in that iteration, it holds that $p^{-}_{i}(\widehat{S}_{i})\leq\overline{\tau}$.
Lemma 31.
Any agent $i\in[n]$ is included at most once in the set $\mathcal{Q}$ throughout the execution of $\mathrm{XEF}1$AugmentTree Algorithm.
The corollary below follows from Lemmas 30 and 31.
Corollary 32.
The $\mathrm{XEF}1$AugmentTree Algorithm (Algorithm 2) always terminates and returns a complete allocation $\widehat{\mathcal{S}}=(\widehat{S}_{1},\ldots,\widehat{S}_{n})\in\mathcal{M}^{*}$.
We now establish Theorem 27.
Proof of Theorem 27.
To prove the theorem it suffices to show that, for any input allocation $\mathcal{S}\in\mathcal{M}^{*}$ with $\varphi^{-}(\mathcal{S})<n$, $\mathrm{XEF}1$AugmentTree Algorithm returns another allocation $\widehat{\mathcal{S}}\in\mathcal{M}^{*}$ with a strictly higher $\varphi^{-}(\cdot)$ value. In particular, by repeatedly applying the algorithm, we obtain the desired allocation $\mathcal{A}^{*}\in\overline{\mathcal{M}}^{*}$ with $\varphi^{-}(\mathcal{A}^{*})=n$, thereby establishing the theorem.
Corollary 32 ensures that the algorithm successfully finds $\widehat{\mathcal{S}}\in\overline{\mathcal{M}}^{*}$. We now complete the proof by showing that $\varphi^{-}(\widehat{\mathcal{S}})>\varphi^{-}(\mathcal{S})$.
Note that, at the end of the first iteration, the root agent $r$ receives a bundle with $p^{-}_{r}(\cdot)$ value at most $\overline{\tau}$ (Lemma 30). Since $r$’s bundle is not modified thereafter (Lemma 31), we have $p^{-}_{r}(\widehat{S}_{r})\leq\overline{\tau}$ in the returned allocation as well.
Moreover, for any agent $j\in[n]$, if in the input allocation $\mathcal{S}$ we have $p^{-}_{j}(S_{j})\leq\overline{\tau}$, then $p^{-}_{j}(\widehat{S}_{j})\leq\overline{\tau}$ in the returned allocation $\widehat{\mathcal{S}}$ as well: If $j$’s bundle was not updated during the algorithm, the assertion holds trivially, since $\widehat{S}_{j}=S_{j}$. Otherwise, if $j$’s bundle was updated and its $p^{-}_{j}(\cdot)$ value became more than $\overline{\tau}$, then, by $j$’s inclusion in $\mathcal{Q}$ and the subsequent updates, the algorithm ensures that $p^{-}_{j}(\widehat{S}_{j})\leq\overline{\tau}$ at termination; recall that $\mathcal{Q}=\emptyset$ when the algorithm terminates.
This observation and the inclusion of $r$ (as an agent which satisfies $p^{-}_{r}(\widehat{S}_{r})\leq\overline{\tau}$) imply that $\varphi^{-}(\widehat{\mathcal{S}})>\varphi^{-}(\mathcal{S})$. The theorem stands proved. ∎
Appendix B Scope for $\mathrm{EF}1$
This section provides an example to show that, starting from Theorem 10, the best one can hope for is an $\mathrm{IEF}1$—and not an $\mathrm{EF}1$—guarantee.
Consider an instance with three agents ($n=3$) and three items, $\{g_{1},g_{2},g_{3}\}$. The valuations of the items (all goods) are listed in Table 1.
$g_{1}$ $g_{2}$ $g_{3}$
$v_{1}(\cdot)$ 30 1 1
$v_{2}(\cdot)$ 30 1 1
$v_{3}(\cdot)$ 30 3 3
Table 1: Valuations of the three agents for the three goods.
As detailed in Section 4.1, we additively perturb the given valuations $v_{i}(t)$s—by random $\varepsilon_{i,t}$s—to construct the values $\widehat{v}_{i}(t)$s in non-degenerate instance. In particular, let $\varepsilon_{1,1}$, $\varepsilon_{2,1}$, and $\varepsilon_{3,1}$ be the perturbations for the first good, $g_{1}$, among the three agents, respectively. Also, let $\eta>0$ be as defined in Equation 4.
Now, consider an weight vector $w^{*}=(w^{*}_{1},w^{*}_{2},w^{*}_{3})$ whose nonnegative components sum up to one and satisfy
$(w^{*}_{1}+\eta)\left(30-\varepsilon_{1,1}\right)=(w^{*}_{2}+\eta)\left(30-\varepsilon_{2,1}\right)=(w^{*}_{3}+\eta)\left(30-\varepsilon_{3,1}\right)$ (42)
Since $\varepsilon_{i,t}$s and $\eta$ are sufficiently small, there exist such $w^{*}_{1},w^{*}_{2},w^{*}_{3}\approx 1/3$.
[…]
$g_{2}$ $g_{3}$
$p^{*}$ $\left(w^{*}_{3}+\eta\right)\left(3-\varepsilon_{3,2}\right)$ $\left(w^{*}_{3}+\eta\right)\left(3-\varepsilon_{3,3}\right)$
We next note that the above-mentioned weight vector $w^{*}\in\Delta_{n-1}$ and optimal prices $p^{*}$ satisfy Theorem 10. Here, in any optimal allocation $(A^{*}_{1},A^{*}_{2},A^{*}_{3})\in\mathrm{OPT}(\textbf{{LP}}(w^{*},\eta))$, goods $g_{2}$ and $g_{3}$ are necessarily assigned to agent $3$. That is, for any optimal allocation $(A^{*}_{1},A^{*}_{2},A^{*}_{3})$ it holds that $\{g_{2},g_{3}\}\subseteq A^{*}_{3}$. The good $g_{1}$ can be allocated to any one of the three agents while maintaining optimality. [^16] The auxiliary good introduced in the non-degenerate instance has a sufficiently low value and, hence, its allocation does not impact the arguments here.
These observations ensure that $w^{*}$ and $p^{*}$ conform to the guarantee in Theorem 10. In particular, each agent $i\in[3]$ can be made price envy-free by assigning $g_{1}$ to it. Specifically, for each $i\in[3]$, including $g_{1}$ in $A^{(i)}_{i}$ gives us an optimal allocation $\mathcal{A}^{(i)}$ wherein $i\in\operatorname*{arg\,max}_{k\in[3]}\ p^{*}(A^{(i)}_{k})$. For instance, for $A^{(1)}_{1}=\{g_{1}\}$ along with $A^{(1)}_{2}$$\emptyset$ and $A^{(1)}_{3}=\{g_{2},g_{3}\}$, we have $p^{*}(A^{(1)}_{1})\geq 9>p^{*}(A^{(1)}_{2})$ and $p^{*}(A^{(1)}_{1})\geq 9>p^{*}(A^{(1)}_{3})$.
Also, observe that this allocation $\mathcal{A}^{(1)}=(A^{(1)}_{1},A^{(1)}_{2},A^{(1)}_{3})$ is $\mathrm{IEF}1$ (under the given valuations $v_{i}$s) – agent $1$ is envy-free under $\mathcal{A}^{(1)}$ and agents $2$ and $3$ can (hypothetically) include good $g_{1}$ into their own bundle to eliminate their existing envy.
However, in this setting, no allocation in $\mathrm{OPT}(\textbf{{LP}}(w^{*},\eta))$ is $\mathrm{EF}1$: As mentioned previously, under any optimal allocation $\mathcal{A}=(A_{1},A_{2},A_{3})$ goods $g_{2}$ and $g_{3}$ are necessarily assigned to agent $3$. Now, either the first or the second agent will not receive $g_{1}$ in $\mathcal{A}$. For such an agent $i\in\{1,2\}$, we have $v_{i}(A_{i})=0<v_{i}(A_{3})$ and the envy persists even after removal of any good from $A_{3}$, since $v_{i}(A_{3}\setminus\{g\})\geq 1$ for each $g\in A_{3}$.
Therefore, starting with a weight vector $w^{*}$ and prices $p^{*}$ as in Theorem 10, it is not necessary that we can identify an $\mathrm{EF}1$ allocation in $\mathrm{OPT}(\textbf{{LP}}(w^{*},\eta))$.